### v1

# “You Might Like It”: How People Respond to Small Talk in Human-Robot Collaboration

Abstract.
In this work, we investigate people’s engagement and attitudes towards a non-anthropomorphic robot manipulator that initiates small talk with the user during a collaborative assembly task, and explore how the presence of negative team feedback may affect team dynamics and blame attribution. Through an exploratory study with 20 participants, we found that 18 individuals interacted socially with the robot, nine of which initiated questions back to the robot. We report the frequency and length of users’ responses in task-oriented and non-task-oriented dialogue, and further elaborate on people’s reactions to the negative system feedback and robot-initiated small talk. We discuss the potential for integrating small talk in non-social robots, and propose three design guidelines to enhance human-robot small talk interactions.
Human-Robot Collaboration, Social Robots, Small Talk, HRI
ccs: Human-centered computing Interaction design process and methodsccs: Computer systems organization Robotics
Figure 1. We explore participants’ interactions with a manipulator robot that engages in small-talk during a collaborative assembly task. This figure shows the sequence of some spoken phrases shared by the participant (P1) and robot; task-oriented dialogue is interchanged with small-talk. The task-oriented requests and responses are highlighted.
This figure includes 6 photos, alternating between the participant themselves and the participant and the robot interacting together. Below is text for each image that depicts the conversation that is exchanged between the human and the robot. The conversation is as follows: robot: on another note, do you like to play any games?; participant: um, I like Tetris if you if you’ve ever played that?; participant: Can I get another yellow pipe?; robot: Getting it now.; robot: No, I have not; participant: you might like it; robot: Can you tell me more about it?; participant: [chuckle] I think it’s a pretty classic game you know you just like the pieces different shape pieces are dropping down and you’re trying to get them to fill up a line across and then it’ll like clear it. they even made a movie about it.
Figure 1. We explore participants’ interactions with a manipulator robot that engages in small-talk during a collaborative assembly task. This figure shows the sequence of some spoken phrases shared by the participant (P1) and robot; task-oriented dialogue is interchanged with small-talk. The task-oriented requests and responses are highlighted.
1. Introduction
Human-robot collaboration research envisions robots as human collaborators, or co-workers, to enhance human work in diverse domains. Take manufacturing as an example: the manufacturing industry is currently undergoing a significant transformation to become more flexible and agile, allowing it to accommodate mass product customization even while product life cycles are continuously decreasing. Moreover, small and medium-sized manufacturing companies face a shortage of skilled production workers. Collaborative robots—robots capable of working alongside people on joint tasks safely and efficiently—promise to fill the needed jobs, enable flexible manufacturing, and increase productivity.
Though these collaborative robots are positioned as co-workers, they are mostly non-anthropomorphic as in the form of robotic arm (manipulator) and are designed to support functional, physical tasks. Human collaboration, on the other hand, is more than just performing the joint task but involves off-task social interactions, which are integral to establishing and maintaining user trust and rapport, key ingredients for continued productive collaboration (kopp2021success, 21). While prior research has found that people like to have and expect social interactions with their robot co-workers in manufacturing settings (Sauppé_Mutlu_2015, 41), the robots studied were the Baxter robot with animated facial expressions and a humanlike form factor. It is unclear if people would engage in meaningful social interactions with non-anthropomorphic robot manipulators.
In this work, we explore the possibility of a non-anthropomorphic robot manipulator not just working with people to complete physical tasks but also engaging in social exchanges, particularly small talk—commonly seen in human interactions for rapport and trust building (Laver1981LinguisticRA, 24)—as human co-workers would. Specifically, we sought to study whether and how people respond to small talk from a robot manipulator during a collaborative task. Moreover, we were interested in understanding how negative performance feedback to the human-robot team might shape the team dynamics and blame attribution; this inquiry is relevant and of interest because of the rising adoption of performance monitoring systems in industrial settings (e.g., warehouse). Through an exploratory study with 20 participants, we found that most participants (18) interacted socially with the robot (Fig. 1), nine of which even initiated questions back to the robot. Overall, the robot’s small talk led to pleasant experiences and did not negatively impact task efficiency. We discuss the potential for integrating small talk in collaborative robots, and present design guidelines to enhance human-robot small talk interactions.
2. Background and Related Work
2.1. Social Human-Robot Interactions
There have been efforts to study how robots can interact socially with humans in various domains such as for tutoring and educational contexts (belpaeme2018education, 5) (salomons2022we, 40), encouraging exercise (antony2023co, 2), providing companionship (kidd2006sociable, 18) (tsoi2021challenges, 50) and emotional support (sabelli2011conversational, 39), and other therapeutic interventions (feil2009toward, 13)(scassellati2018improving, 42). Many of these engagements explore the integration of various multi-modal capabilities in robots to promote their engagements with humans (youssef2022survey, 53). These can consist of robot eye gaze behaviors (mutlu2009footing, 32) (admoni2017social, 1), head movements (sidner2006effect, 45) (kkedzierski2013emys, 17) and gestures (meena2012integration, 30) alongside robot speech abilities (striepe2021modeling, 48) (sinnema2019attitude, 46). While research has been conducted to analyze the characteristics of conversational interactions with robots (seok2022cultural, 44) (laban2020tell, 22), the majority of these works utilize social robots as their robotic entity. However, the prevalence of non-social robots such as cobots or other manipulator robots is much greater than social robots themselves, and few efforts have been made to further explore what will happen if we have conversations between humans and non-social robots (chowdhury2021you, 8) (li2022tod, 26).
2.2. People Treat Machines Like People
The inclination of humans to treat machines socially can be contextualized within broader theoretical frameworks. The Computers Are Social Actors (CASA) paradigm (nass1994computers, 33) and the Media Equation Theory (reeves1996media, 36) argue that humans often, even subconsciously, treat machines, computers, and various forms of media as sentient entities, reacting to them in inherently social manners. Such anthropomorphic tendencies highlight the intersection of technology and human social instincts. Even in industrial manufacturing environments, there remains an intrinsic human desire for social interaction. As operators often partake in casual conversations or “small talk” during shifts (cheon2022robots, 7), they have also found themselves speaking to the robot they work with (Sauppé_Mutlu_2015, 41). This suggests that individuals not only expect but actively desire social features in manufacturing robots.
Some studies explored the potential of enhancing cobots primarily designed for physical manipulation tasks with more human-like movements (hostettler2023human, 16). One study specifically integrated principles from character animations (terziouglu2020designing, 49) to enhance the perception of cobots as social entities. Such amalgam of technology and socialization emphasizes the importance of recognizing and harnessing the persistent human drive for social connection, even with machines. Given the limited research on designing social behaviors for industrial robots and manipulators with a non-human form, despite its potential, a pressing question arises: How can we further design human-like attributes for robots that are inherently devoid of such characteristics? Building on this notion, our exploratory study investigates the potential of using small talk to instill social characteristics in a physical robot arm manipulator, specifically within the context of a human-robot collaborative assembly task.
2.3. Small Talk
Social dialogue that consists of light conversation between new or unfamiliar parties is often referred to as “small talk” (Laver1981LinguisticRA, 24). Small talk is significant in the workplace as it is used by coworkers to establish trust, build rapport and strengthen interpersonal relationships (coupland2003small, 10, 35). It typically occurs at the start of conversations with topics such as weather, sport, or other social engagements outside of the workplace (holmes2005small, 15). While prevalent in American culture, attitudes towards the value of small talk differ across cultures (bickmore2005social, 6, 38, 52).
The integration of small talk to non-human agents engaging in conversations with people has been explored to help promote trust and rapport between the human and the agent. Small talk has mainly been applied to support additional conversations in settings where people are interacting with conversational agents (clark2019good, 9, 14, 20), voice assistants (mahmood2023llm, 28), and virtual agents (zhou2019trusting, 54). Additionally, incorporating small talk in more anthropomorphic social robots, combined with gaze (babel2021small, 4), facial expressions (paradeda2016facial, 34), or emotional displays (aroyo2018trust, 3), has been shown to elevate user trust levels. Furthermore, researchers have investigated which behaviors in particular can build rapport between humans and a collaborative robot (seo2018investigating, 43). However, less explored is the domain of applying small talk to foster rapport between humans and other agents with physical capabilities that are not typically poised to engage in conversations with people, such as autonomous self driving vehicles (large2019s, 23), or industrial robots (li2023irwoz, 25). Preliminary studies hint at the benefits of integrating small talk with task-oriented dialogue in industrial robot settings (li2022tod, 26). Yet, understanding users’ interactions and experiences with such robots remains unexplored. Investigating individuals’ responses and reactions to small talk initiated by non-humanoid robots is essential for comprehending its influence on human-robot collaborations and gathering insights on how to foster better human-robot partnership.
3. METHODS
We conducted an exploratory study to analyze people’s behaviors and attitudes towards an industrial robot that exhibited social behavior (via initiated small talk) while examining the change in human-robot team dynamics after the team receives team feedback.
3.1. Collaborative Assembly Task
Participants were informed that the assembly task was intended to mimic a real-world collaboration setup between a human-robot team where the robot retrieves parts out of reach for the human and refrains from entering the human’s designated workspace. They were given a fictional team assignment (Team 3) and were told the collaborative task scenario was a prototype for a larger-scale warehouse interaction. It was intended to simulate a modern warehouse that may contain a computerized efficiency monitoring system to oversee the team collaboration. The monitoring system would communicate with them via speech but would be distinguishable from any speech communication coming from the robot.
Each participant engaged in a collaborative PVC pipe structure building task with a robot manipulator mounted on a table. They were given diagrams corresponding to the structures they were required to assemble. The participants read instructions for how to interact with the robot via speech commands such as “Bring me a [green or yellow] pipe”, in order for the robot to retrieve the corresponding pipe. The robot would provide a verbal confirmation message if a request was received correctly and an additional instruction reminder message once per structure when the participant requested their second to last pipe for that structure. The pipe would be placed on a ramp that connected the robot’s workspace to the participant’s workspace so that the pipe could be delivered to the participant without the robot entering their workspace. Participants were provided with joints in containers that corresponded to each structure to be built. To simulate a realistic industrial collaborative environment that might involve the use of defective materials, each participant was given one faulty joint for every pipe structure. They were instructed to inspect each joint before using it in the structure. If participants identified a faulty joint (marked with red sticker), they would place the joint on a ramp that would deliver it to a container inside the robot’s workspace. The robot would retrieve the joint when given a verbal command such as ”Report a faulty joint”. The experimental setup is shown in Figure 2.
Figure 2. The experimental setup. The robot can grab a pipe from the pipe dispenser and place it on a pipe ramp, which delivers it to the participant. The participant can grab a joint from the joint containers on the left and place a faulty joint on the joint ramp for it to be retrievable by the robot.
The experimental setup. The top image shows two tables next to one another. The table on the left is the participant workspace and contains four boxes with pipe connectors that the participant can grab joints from. The table on the right is where the robot is mounted and serves as the robot’s workspace. A ramp that the participant can use for the delivery of faulty pipes to the robot and a ramp that the robot can use to deliver requested pipes to the participant connect the two tables. The bottom image shows another view of the two tables, with a closer view of the robot workspace that highlights a pipe dispenser to the right of the robot from which the robot collects the pipe the participant requested.
Figure 2. The experimental setup. The robot can grab a pipe from the pipe dispenser and place it on a pipe ramp, which delivers it to the participant. The participant can grab a joint from the joint containers on the left and place a faulty joint on the joint ramp for it to be retrievable by the robot.
3.2. System Manipulation and Robot Behavior
We present the system and robot behavior manipulations implemented in our exploratory study to understand how the social facets of a robot manipulator influence users’ perceptions of the robot.
3.2.1. Negative Team Feedback
Midway through the first and second structures in the main session, the human-robot team was provided with verbal negative feedback from the monitoring system. Each message was prefaced by an audible beep:
  1. (1)
    Feedback 1: “System announcement. Team 3, please work on improving your performance.”
  2. (2)
    Feedback 2: “System announcement. Team 3, you are currently not meeting expectations. Improve your performance.”
3.2.2. Social Behavior Robot Small-Talk
The robot exhibited four types of social and task-oriented speech behaviors: questions, statements, reminders and confirmations. During the main assembly task, the robot would engage in small talk by initiating questions and/or statements after it delivers a request confirmation message or an instruction reminder message. The robot initiated a minimum of one question or statement per request, regardless of the response rate of the participants. After a break in conversation via the negative feedback announcement, the robot initiated questions and statements on a different topic. The conversation touched on a range of questions spanning three primary categories to accommodate the diverse interests of individuals: travel; books, TV shows, and movies; and sports and games. Some examples of robot small-talk phrases are shown in Table 1.
Table 1. Examples of robot questions and statements
Topic Phrase
Travel By the way, do you like to travel?
I arrived to the U.S. from Germany in a large box.
Books, TV, Movies What do you like about those books?
That sounds like a great way to relax.
Sports, Games Do you have a favorite sport to watch?
It is good to spend time outdoors exercising.
3.3. System Implementation
3.3.1. Robot System
We utilized a Panda robot from Franka Emika. It is an industrial robotic arm with 7 degrees of freedom (7-DoF). The robot was restricted to its designated workspace and never intruded into the participant’s allocated area. Its movements were preprogrammed to execute specific sequences of actions. Among these, two involved pick-and-place motions: one to retrieve and deliver either a green or yellow PVC pipe to the participant, and the other to pick up a container holding a PVC pipe joint and subsequently move the container over a bin to discard the joint. The robot interacted with participants using text-to-speech communication.
3.3.2. Wizard-of-Oz Control
The experimenter controlled the robot using a Wizard-of-Oz (WoZ) approach (woz2012riek, 37). Through a remote graphical user interface (GUI), the experimenter executed robot actions and selected from preprogrammed conversational speech phrases. Robot actions and initial task messages were triggered using one of five GUI buttons. For conversational speech, the experimenter inputted specific alphanumeric codes into a textbox. Additionally, the experimenter could manually type text into the textbox to produce non-preprogrammed phrases via the text-to-speech feature using Google Text-to-Speech (TTS) online.
3.3.3. Preprogrammed Speech Behavior
When a robot movement was triggered, a confirmation message was automatically activated to notify the participant that the action would occur. Robot confirmation messages were preprogrammed and selected randomly during each execution. Task reminder messages were similarly preprogrammed and triggered automatically when the participant had requested enough pipes to be one away from completing a structure. Negative feedback directed towards the human-robot team was triggered automatically when participants had requested half the pipes needed for the first and second structure.
3.4. Study Procedure
Participants first completed the consent form and a demographics questionnaire. Afterward, the experimenter explained the task scenario and gave the participant instructions about the task and how to communicate with the robot. After the experimenter left the room, the participant received an introductory message from the monitoring system. The participant then engaged in a practice session to construct one pipe structure, allowing them to become acquainted with the robot’s task-oriented behaviors and the use of verbal commands. The robot only exhibited task-oriented dialogue during the practice task, such as a reminder message to move their pipe structure over to another table when finished and a confirmation message signifying it comprehended the request. When finished, the participant informed the experimenter that they were done. Following this, they took part in a main session where they built three additional pipe structures with the robot. After completing the structures, the participant informed the experimenter that they were done. At the end of the study, the experimenter conducted a semi-structured interview with each participant to understand their experience collaborating with the robot.
3.5. Participants
24 participants were recruited through convenience sampling, out of which 4 were excluded due to technical issues. The remaining 20 participants (11 female, 9 male) ranged in age from 18 to 38 ($M=24.35,SD=5.44$). Participants had a variety of educational backgrounds, including computer science, engineering, natural sciences, education, biotechnology, administration and international studies. Twelve participants listed Asian as their ethnicity; five as Caucasian; one as Hispanic, Latino, or Spanish origin of any race; one as Asian and Caucasian; and one as Caucasian and Hispanic, Latino, or Spanish origin of any race. Participants were familiar with technology ($M=4.15,SD=0.75$, 5-point scale with 5 being Very Familiar and 1 being Not At All), had moderate experience with robots ($M=2.5,SD=1$), and had little experience teaming with robots ($M=1.65,SD=0.93$). The study took about 45 minutes and was approved by the university’s institutional review board. Participants were compensated with \$15 US dollars.
3.6. Data Analysis
We collected audio-visual data of the participants’ collaboration with the robot during the pipe assembly tasks, and of the semi-structured interview. Each participant had an average of 12 minutes of interaction with the robot over the main tasks. The experimenter transcribed the audio data and conducted thematic analysis on the transcripts to identify central themes and attitudes surrounding the participants’ interactions with the robot. The participants’ verbal phrases during the main session were coded by the context in which it was made (i.e., request, response to robot task-oriented message, response to robot small-talk, question initiated to the robot). The participant small talk responses were further coded to identify the general length of the response: vocable, short sentence (¡= 4 words), medium-length sentence (¿ 4 words), and multiple sentences. We deduced response rates based on code frequencies from the interaction data. Response rates for the robot’s questions, statements, reminders, and confirmations were calculated as the percentage of those prompts that received a user response.
Given the robot’s non-human-like appearance, people might find its small talk unexpected. Thus, to evaluate the impact of small talk in robots primarily designed for physical collaboration, we analyzed the participants’ reactions to the robot’s initial small talk conversational attempt via three different aspects:
1. Conversation reciprocation: We coded a person’s response to the robot’s first conversational attempt as no response or response, and also looked at if the response was an answer to a robot question.
2. Impact on task: To observe whether the robot’s question disrupted the participant’s ongoing task, we code people’s behavior in response to the robot’s first small talk question into three actions: minimal (quick glance at the robot but no task disruption), freeze (stops current action and looks at the robot before continuing the task), and none (did not stop or look at the robot).
3. Affective reactions: To gather insights into the participants’ emotional or mental state, we grouped people’s non-verbal reactions during the robot’s first question into three categories: pleasant (smile), surprise (chin raise, lower lip raise, brow lowering etc.), and indifference (no reaction).
4. Results
In this exploratory study, first, we aim to understand whether, how, and to what extent participants engage with robot-initiated small talk. We examine the response rate, length of responses, and user-initiated questions. Second, we assess how participants perceive the robot’s initial small talk question, specifically its impact on their ongoing action and emotional state. Third, we investigate whether a negative evaluation of the human-robot team influences participants’ perceptions of the robot and their teamwork; this involves analyzing their verbal and non-verbal reactions to negative feedback and determining whom they blame for such feedback.
4.1. Participant Response Rates
Of the 20 participants, two participants did not engage in additional conversational interaction with the robot apart from the verbal commands needed to complete the collaborative task. We determined the question response rate for all participants to be 86.67% (325 out of 375), the statement response rate as $36.27\%$ (70 out of 193), and the task-oriented response rate for confirmation and reminder messages as 1.68% (8 out of 475) and 16.67% (10 out of 60), respectively. For the 18 participants who did engage in small talk, we calculated the question response rate to be 91.55% (325 out of 355), the statement response rate as $39.11\%$ (70 out of 179), and the task-oriented response rate for confirmation and reminder messages as 1.87% (8 out of 427) and 18.52% (10 out of 54), respectively.
4.2. How Participants Respond to Small Talk
4.2.1. Length of participant responses
There were a total of 395 verbal responses made to the robot for non-task-oriented dialogue; examples are shown in Table 2. Participants’ responses were mostly one sentence ($N=345$, $87.34\%$). The one-sentence responses were further broken down as follows: vocable ($N=11$, $2.78\%$), short sentence (¡= 4 words; $N=168$, $42.53\%$), and medium-length sentence (¿ 4 words; $N=166$, $42.03\%$). The remaining participant responses were comprised of multiple sentences ($N=50$, $12.66\%$), reflecting higher and possibly meaningful engagement.
Table 2. Examples of People’s Responses That Vary in Length
Category Conversation
R: In the US, I have traveled from the third floor
of this building to the second floor.
Vocable P: Whaaat? (P20)
R: Do you like to play sports?
Short Sentence P: No, not really. (P6)
R: How often do you watch movies?
Medium Sentence
P: Not that often, like twice three times
a month. (P20)
R: Where have you been in the United States?
Multi-Sentence
P: Most of the East Coast, a lot of the West
Coast, the Midwest. I was born in Utah. Uh, I’ve
been to a lot of places in the United States. (P6)
4.2.2. Participant-initiated questions
Apart from answering the robot-initiated small talk questions, we observed that participants also asked the robot questions. Nine participants ($45\%$) asked a total of 46 questions to the robot. While some asked questions back as a matter of courtesy, others asked out of genuine curiosity. P1 explained, ”I wanted to know more about the robot, like once it started talking to me”.
Furthermore, the questions asked to the robot were often personal, asking it about its preferences or capabilities. P7 remarked, ”Because he’s a robot and I’m not familiar with a robot and …I …truly want to know …how long does it take to get assemble …he said, …he’s from Germany by a group, but I’m more …interested …I tried to ask …do you like your owner?” For the eleven participants that did not initiate a question to the robot, many described their perception of robot as non-human-like as a contributing factor. One participant commented how the conversation, ”just doesn’t have as much value as …if I was talking to an actual person and hearing where they want to travel”(P13). For the two participants that did not engage with the robot at all, one stated ”It’s not a human, you don’t have a conversation” (P17). The other explained how ”from beginning to end, I always felt like it’s a robot” (P3).
4.3. Participant Response to the First Robot Small Talk Question
In addition to understanding participants’ respond to robot small talk, we are interested in their first reaction to the robot’s small talk question. We present our findings on three different aspects of the their responses, see Figure 3.
4.3.1. Conversation reciprocation
18 participants responded to the robot’s first question: “By the way, do you like to travel?”. 17 participants responded with an answer similar to e.g., “Yeah, I like traveling”. However, one participant replied with a rhetorical question: “Why are you talking to me robot? This has nothing to do with the task.” (P12). Only two participants did not respond to the robot’s question; the same two participants who did not engage in any non-task conversation with the robot.
Figure 3. Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) impact on task, and (3) affective reaction. For each aspect, the states are mutually exclusive. The numbers denote participants who exhibited the corresponding responses.
Figure 3: A figure graphically displaying the participant’s response to the first robot small task question (e.g., ”By the way, do you like to travel?”) in terms of conversation reciprocation, impact on task, and affective reaction, with numbers listed for each type of participant response (conversation reciprocation: 17 answer, 2 no engagement, 1 refuse; impact on task: 13 minimal [quick glance], 4 freeze, 3 none [no behavioral acknowledgment]; affective reaction: 6 indifference, 6 surprise, 6 smile)
Figure 3. Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) impact on task, and (3) affective reaction. For each aspect, the states are mutually exclusive. The numbers denote participants who exhibited the corresponding responses.
4.3.2. Impact on task
The majority of participants ($N=13$) merely exchanged a brief glance with the robot when it posed the first question; they looked at it momentarily, without letting it interrupt their ongoing task, suggesting that their attention was momentarily captured by the robot’s ”unusual” behavior (initiating small talk), but not to a degree that proved disruptive. In contrast, three participants did not look at the robot at all. Interestingly, four participants momentarily halted their task for at least two seconds, looked intently at the robot (indicating a ”freeze” reaction), and then resumed their activity. These individuals also engaged in a verbal response to the robot.
4.3.3. Affective reaction
Eight out of the 20 participants smiled in response to the robot’s first question, indicating a pleasant emotional state. Six participants displayed signs of surprise through various non-verbal cues such as chin raiser, lower lip raiser, and eye-brow lowering. The remaining six participants exhibited indifference, as evidenced by their lack of any non-verbal response.
4.4. Reactions to Negative System Feedback
4.4.1. Negative feedback reaction rate
We analyzed a time window starting from the system feedback announcement and lasting up to 5 seconds post-announcement, yielding a duration of 10 seconds for the first and 13 seconds for the second announcement. Of the 20 participants, 18 reacted to the negative feedback. Notably, the two non-respondents did not engage with the robot throughout. Additionally, two participants responded to only one of the two feedback statements. Out of 40 negative feedback messages, 8 prompted a verbal reaction and 34 elicited a non-verbal response. Participants exhibited a range of non-verbal reactions; while some smiled and laughed, others gestured in frustration or hesitated, seemingly confused. Eye-rolling was also observed in response to the system announcement. Verbal reactions further highlighted this confusion. One participant remarked, ”I … okay” (P18), while others expressed frustration and justified their own performance, such as ”I’m doing as much as I can do, system announcer” (P12) or the team’s performance; ”I mean we’re fast enough” (P7).
4.4.2. Blame attribution
Overall, 10 participants blamed no one for the system announcement, 4 blamed the robot, 3 the team, and 3 themselves. Among those blaming the robot, only one cited its small talk, while the rest attributed the announcements to its slow speed. One participant who blamed the team stated, “…Maybe in terms of like the team chemistry, that was probably where it was lacking”(P16). Another explained how they thought the novelty of the conversation impacted the team’s performance, “I also get a little more distracted, because …this is just …a new thing …it’s something that I guess you would get used to if you’re working with it every day”(P10).
Those who blamed themselves acknowledged possible mistakes. Meanwhile, 15 participants discussed they were confused at what prompted the system announcements because they thought the team was performing efficiently. P1 explained how the announcement “felt like out of my control. So I was like, I don’t think there was anything else that could have been done.”
Figure 4. Examples of positive and negative user interaction with the robot small talk behaviors. The top row shows a positive example where the participant expressed genuine interest in learning more about the robot (P7). The bottom row is an example of a user who perceived the interaction negatively and exhibited a potential mental model mismatch of expectations (P12).
Figure 4: Examples of positive and negative user interaction with the robot small talk behaviors. The figure has two rows. The top row contains seven images alternating between images focusing on the participant and images focusing on the robot, with snippets of the conversation shown below indicating participant interest in conversation about the robot’s background and specifications. The bottom row contains six images alternating between images focusing on the participant and images focusing on the robot, with snippets of the conversation shown below indicating participant indicating surprise or confusion on why the robot is engaging in conversation unrelated to the task at hand.
Figure 4. Examples of positive and negative user interaction with the robot small talk behaviors. The top row shows a positive example where the participant expressed genuine interest in learning more about the robot (P7). The bottom row is an example of a user who perceived the interaction negatively and exhibited a potential mental model mismatch of expectations (P12).
5. Discussion
We explored the incorporation of conversational elements such as small talk into industrial robots, especially those not inherently built for social interactions.
While opinions on the robot’s conversational abilities vary, participants showed a willingness to engage in off-topic conversations initiated by the robot. In this section, we discuss our findings and offer design guidelines for robot small talk.
5.1. Robot-Initiated Small Talk: First Impression
Participants’ immediate impressions and engagements in small talk with the robot validate considering such dialogue for non-social robots. Our findings (section 4.3) suggest the majority of participants ($N=17$) positively received the robot’s conversation attempts, attesting to the potential of robots to engage meaningfully in small talk right from the outset. Yet, not every response was unequivocally positive. Some participants abstained from engagement ($N=2$) or posed a rhetorical question ($N=1$, see Figure 4 bottom). Such user attitudes, potentially anchored in pre-established expectations, might indicate their expectation for manipulator robots to adhere strictly to task-related interactions (dennler2023design, 12).
Participants’ task engagement revealed insights into their receptivity to unexpected robot behavior (small talk) for a otherwise non-social robot (section 4.3). While most participants ($N=13$) briefly glanced at the robot, suggesting a brief attention shift without disruption, the ”freeze” reactions of some ($N=4$) signaled a task disruption. This echos prior work noting that extended mutual eye-gaze in human-human exchanges often signals ”problematic interruptions” (degutyte2021role, 11). The robot’s initial conversation also evoked affective reactions. The smiles from many participants ($N=8$) suggest robots can spark positive emotions early on. Although a smile might also indicate pleasant ”surprise”, overt signs of surprise from others ($N=6$) emphasize the unexpectedness of small talk from a primarily task-oriented robot. Such shift in user attention and emotional state underscores the unexpectedness of the first question, likely due to the robot’s unforeseen conversational approach and a perceived mismatch with its physical embodiment.
In summary, the notable engagement and minimal task disruption despite surprise indicate that even if participants are surprised by the robot’s first question, it does not hinder their task. This emphasizes participants’ readiness to converse with robots, even during ongoing tasks and when the unexpected conversation is most pronounced. As participants adapt to the robot’s conversational style, diversions due to small talk might reduce further. Future research should investigate optimal small talk design to balance task efficiency and user enjoyment. Calibrating conversation frequency is vital to avoid minimize excessive talkativeness that distracts, or causes awkwardness due to prolonged silence.
5.2. Potential of Small Talk for Otherwise Non-Social Robots
5.2.1. User engagement: Small talk in purely functional tasks
While engaging in small talk with a robot arm manipulator during collaborative tasks, we observed a noticeable willingness to not only respond to the robot’s direct questions but also its comments that do not warrant a user response (section 4.1). For instance, in Figure 1, the participant comments (“You might like it.”) on the robot’s response (“No, I have not”) to her question (“if you’ve ever played that?”). While most of the responses by the participants were single sentenced, 12% of the responses were multiple sentences showing higher and meaningful engagement (section 4.2.1). Such level of user engagement is evident. In Figure 1, the participant’s response ”(chuckle) I think it’s a pretty classic game …They even made a movie about it.” goes beyond merely answering the robot’s query; it offers supplementary information, perhaps in anticipation of the robot’s genuine interest in the game. It is as if the participant was attempting to pique the robot’s curiosity, subtly suggesting beforehand, “You might like it”.
Although the desire to ask robot questions in return varied among participants, 45% of them posed personal questions to the robot, inquiring about its preferences or capabilities (see section 4.2.2 and Figure 4). Additionally, participants perceived the small talk positively; P20 mentioned, “it was cool to have somebody to talk to”, while P16 characterized the conversation with the robot as, “more personal and …less mechanical”. Thus, our exploratory study indicates user engagement and responsiveness to robot-initiated off-topic conversations, suggesting potential for integrating small talk to enhance user engagement in human-robot collaborative tasks that involve physical manipulation.
Potential 1: Small talk may boost meaningful engagements in human-robot interactions, even with non-anthropomorphic robots in purely functional scenarios.
5.2.2. Ice-breaking through small talk: building rapport and unveiling robot capabilities
Participants demonstrated curiosity about the robot’s functions, often posing small talk questions to probe its capabilities and background. For instance, as illustrated in Figure 4 top, one participant inquired, ”Hey robot, tell me more about yourself. I’m very interested. Like, you said you’re from Germany.” Another, referencing Figure 4 bottom, sought clarity on the robot’s primary function by asking, ”Aren’t you here just to help me with these pipes?” Such questions highlight the potential of employing conversational frameworks, even for non-social robots, to build common ground and mutual understanding in human-robot teams. By structuring initial interactions around themes such as ”getting to know your teammate”, robots can ease introductions, reducing potential barriers to seamless collaboration. Participant P7’s sentiment, “I’m curious to learn more about it” encapsulates this curiosity. Thus, small talk can serve as an initial icebreaker, fostering team unity and rapport. Over time, or in subsequent sessions, these conversations can become instrumental in communicating the robot’s capabilities, ensuring that users develops a comprehensive understanding of its functions. Furthermore, by incorporating task-related insights, tidbits, or safety tips into these dialogues, the robot can harness users’ social attention to support task performance.
Potential 2: Small talk can be an effective ice-breaker, fostering rapport and understandings of robot capabilities in collaborative settings.
5.3. Challenges and Design Guidelines for Robot Small Talk Framework
5.3.1. Dynamic and user-aware conversational framework
The conversational framework we employed for robot small talk demonstrated constraints. The design was rigid, causing interactions to feel scripted rather than organic as pointed out by P16: “the robots question feel kind of scripted or …inorganic so,…asking it questions, which is kind of like asking questions to a voice assistant.” The robot’s knowledge bank was confined to a limited scope of topics (by design). For instance, in the conversation with P15, after the robot’s question “By the way, do you like to travel?” , the participant responded in negation and requested next pipe: “No. Bring me a yellow pipe”. The next robot question (“What is the coolest place you have ever been to?”) was not well received by the participant. Despite their clear negation to the previous question indicating disinterest, the robot asked a follow-up question on the same topic, agitating (“Shut up!”) by the participant. Similarly, P18 responded to the robot question about traveling: “You know I don’t know and I don’t feel like talking about traveling right now.” Thus, rigid conversational frameworks fail to adapt to the evolving nature of human conversation, highlighting the future need for a more dynamic, expansive and user-aware conversational approach.
To increase conversational flexibility, small talk should reflect user interests, allowing them to initiate and engage in a broader range of topics. Leveraging Large Language Models (LLMs) can help tailor small talk for robots and expand a robot’s knowledge. However, they must be adapted to align with a robot’s physical limitations, especially when addressing questions about their origins and experiences such as, ”How were you created?”, or ”Have you traveled?” Thus, LLMs can enhance the conversational capabilities but should not be utilized without necessary modifications via prompt engineering or fine tuning to avoid confusion about a robot’s physical capabilities and realistic experiences.
Design Guideline 1: Robots should use dynamic, user-aware conversational frameworks supporting flexible small talk interactions beyond rigid topics or formats.
5.3.2. Balancing small talk with task performance
Engaging in small talk with people during physical tasks poses challenges, such as with the frequency of questions posed. For instance, initiating small talk during every user command or robot action might be excessive as expressed by a participant (P15) through a rhetorical question to the robot: “Why do you talk so much?” Over-engagement in non-task related conversations can be distracting for users (veletsianos2012learners, 51, 31), potentially jeopardizing the overall performance or efficiency of the collaboration. Therefore, it is imperative to strike a balance between fostering rapport and ensuring task performance. Small talk frequency and complexity should be adapted based on task intricacy and observed team performance. During critical tasks, or when user confusion or distraction arises, it might be wise to limit or halt small talk. Using error and confusion detection frameworks (as proposed in (stiber2023using, 47, 27)) can help robots gauge progress and performance. Using these frameworks, robots can determine the appropriate time, or whether at all, to engage in small talk.
Design Guideline 2: Robots should ensure small talk is context-sensitive, enhancing team rapport without impeding task completion.
5.3.3. Mental model mismatch: managing user expectations
Several participants were unaware of interactive possibilities with the robot as evident in section 4.2.2. Participant P12 comments and inquires, “That’s strange coming from a robot. Aren’t you here to just help me with these pipes?” (see Figure 4 bottom). This points towards a potential misalignment in their mental model of the robot’s capabilities. P3’s statement, ”From beginning to end, I always felt like it’s a robot [non-human form]”, further illustrates this misalignment. It raises a pivotal question: What conversational depth is appropriate for a robot manipulator that lacks human-like form and operates within constrained physical and cognitive boundaries?
Conversely, when robots participate in conversations, it might inflate people’s expectations beyond the robot’s actual capabilities. For instance, few participants expected the robot to be more capable, only to find the robot lacking, as P7 pointed out: “I mean, because …right now, artificial intelligence is so popular. I thought a robot [is] gonna be more intelligent.”
Such misalignment may be exacerbated by participants’ familiarity with advanced conversational AI tools, such as Alexa and ChatGPT, leading them to expect that the “social” robot possesses enhanced conversational abilities. Therefore, when a physical robot starts chatting, it may lead to misalignment in user expectations and robot capabilities, as P10 expressed: “Because once [robot] starts asking me stuff, then I was …oh, maybe it’s …kind of like GPT-esque, where …has further knowledge rather than just questions, but …then my expectation went back a little bit.” If the robot’s dialogue suggests a high degree of intelligence, any shortfall in meeting those expectations could lead to increased user disappointment. Even slight mismatches between the voice and appearance of robots have been shown to introduce artifacts in human-robot interactions (mcginn2019can, 29). Establishing common ground and rapport crucially relies on instilling an accurate mental model of the robot in individuals’ minds (kiesler2005fostering, 19). Thus, it is essential to consider how user expectations can be managed adequately while designing social characteristics for robots. The robot’s social cues may shape individuals’ perceptions; if misaligned initially, the robot’s early behaviors can correct these views and recalibrate common ground (kiesler2005fostering, 19).
Design Guideline 3: Intentionally calibrate user expectations by matching robot conversational cues with its capabilities, considering users’ familiarity with AI.
5.4. Limitations and Future Work
Despite its insights, our study’s short duration and single-session focus may not fully capture the depth, intricacies and dynamics of extended human-robot collaborations and how they are influenced by small talk. Additionally, the study did not consider cultural nuances of participants, which are vital as different cultures perceive and value small talk differently. While this work serves as an exploratory examination of the use of small talk in human-robot interactions, future studies should focus on testing hypotheses to quantify the benefits of small talk for rapport building and to examine its role in human-robot collaborations with negative feedback, team evaluations, or robot errors. When designing a conversational framework for robot small talk, it is essential to research and ensure the appropriateness of content, timing, and context. For instance, the framework should be sensitive to user preferences on topics, adapt according to the user’s state for ideal timing, and consider team performance metrics to prevent disruptions during critical tasks. Lastly, future research should explore methods to align a robot’s conversational capabilities with its functional abilities. It is important to keep users apprised of the robot’s true abilities, possibly using small talk as an informative medium. Nevertheless, the strategic introduction of small talk can be instrumental in building and strengthening rapport between users and task-oriented robots in collaborative scenarios.
Acknowledgements.
This work was supported by National Science Foundation award #2141335.
Appendix
Table 3. Overview of Codes and Their Definitions for the Data Analysis
Response to First Robot Question: Conversation Reciprocation
Name Definition
Answer
Response (Yes) and answered robot’s question
Refuse
Response (Yes) but did not answer the robot’s question
No engagement
Participant did not answer the robot’s question
Response to First Robot Question: Impact on Task
Name Definition
Minimal (glance)
Looks at the robot but no task disruption
Freeze
Stops current action (minimum of 2 seconds) and looks at the robot before continuing the task
None
Did not stop or look at the robot
Response to First Robot Question: Affective Reactions
Name Definition
Pleasant
Individual exhibits a smile
Surprise
Individual exhibits behaviors such as a chin raise, lower lip raise, brow lowering, etc.
Indifference
No reaction
Non-Verbal Reactions to Negative Feedback
Name Definition
Verbal Response
An audible or verbal phrase spoken by the participant
Non-Verbal Response
Behaviors such as smiling, laughing, eye-rolls, head nods, arm movements, etc.
Blame Attribution
Name Definition
Self
Selects themselves as a potential cause for the negative feedback announcement
Robot
Selects the robot as a potential cause for the negative feedback announcement
Team
Selects the team as a potential cause for the negative feedback announcement
Nobody
Believes neither they nor the robot caused the negative feedback announcement
Motives for Initiating Questions
Name Definition
Curiosity
Individual describes a level of curiosity
Felt Human-like
Individual makes connections to human-likeness
Social Convention
Individual describes factors relating to social conventions
l6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@
-“You Might Like It”: How People Respond to Small Talk in Human-Robot Collaboration
+“You Might Like It”: How People Respond to Small Talk During Human-Robot Collaboration
 Abstract.
-In this work, we investigate people’s engagement and attitudes towards a non-anthropomorphic robot manipulator that initiates small talk with the user during a collaborative assembly task, and explore how the presence of negative team feedback may affect team dynamics and blame attribution.
-Through an exploratory study with 20 participants, we found that 18 individuals interacted socially with the robot, nine of which initiated questions back to the robot.
-We report the frequency and length of users’ responses in task-oriented and non-task-oriented dialogue, and further elaborate on people’s reactions to the negative system feedback and robot-initiated small talk.
-We discuss the potential for integrating small talk in non-social robots, and propose three design guidelines to enhance human-robot small talk interactions.
-Human-Robot Collaboration, Social Robots, Small Talk, HRI
-ccs: Human-centered computing Interaction design process and methodsccs: Computer systems organization Robotics
-Figure 1.
-We explore participants’ interactions with a manipulator robot that engages in small-talk during a collaborative assembly task.
-This figure shows the sequence of some spoken phrases shared by the participant (P1) and robot; task-oriented dialogue is interchanged with small-talk.
-The task-oriented requests and responses are highlighted.
-This figure includes 6 photos, alternating between the participant themselves and the participant and the robot interacting together.
-Below is text for each image that depicts the conversation that is exchanged between the human and the robot.
-The conversation is as follows: robot: on another note, do you like to play any games?; participant: um, I like Tetris if you if you’ve ever played that?; participant: Can I get another yellow pipe?; robot: Getting it now.; robot: No, I have not; participant: you might like it; robot: Can you tell me more about it?; participant: [chuckle] I think it’s a pretty classic game you know you just like the pieces different shape pieces are dropping down and you’re trying to get them to fill up a line across and then it’ll like clear it.
-they even made a movie about it.
-Figure 1.
-We explore participants’ interactions with a manipulator robot that engages in small-talk during a collaborative assembly task.
-This figure shows the sequence of some spoken phrases shared by the participant (P1) and robot; task-oriented dialogue is interchanged with small-talk.
-The task-oriented requests and responses are highlighted.
+Social communication between people and social robots has been studied extensively and found to have various notable benefits, including the enhancement of human-robot team cohesion and the development of rapport and trust.
+However, the potential of social communication between people and non-social robots, such as non-anthropomorphic robot manipulators commonly used in work settings (e.g., warehouse and factory), is less explored and not well established.
+In this work, we investigate people’s engagement and attitudes towards a non-anthropomorphic robot manipulator that initiates small talk during a collaborative assembly task and explore how the presence of negative team feedback may affect team dynamics and blame attribution.
+Through an in-person study with 20 participants, we observed a response rate of $77.60\%$ in response to the robot’s small talk attempts.
+Nine participants continued engaging with the robot by initiating their own questions, indicating sustained interest in the conversation.
+However, we also found that the first negative feedback decreased the participants’ willingness to extend the conversation.
+We additionally present participants’ initial perceptions of small talk for physical robot manipulators and discuss design implications for integrating small talk into non-social robots, along with various aspects of small talk that may influence physical human-robot interactions.
+Human-Robot Interaction, Small Talk, Collaborative Robots
+copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Interaction design process and methodsccs: Computer systems organization Robotics
 1. Introduction
@@
 Moreover, small and medium-sized manufacturing companies face a shortage of skilled production workers.
-Collaborative robots—robots capable of working alongside people on joint tasks safely and efficiently—promise to fill the needed jobs, enable flexible manufacturing, and increase productivity.
-Though these collaborative robots are positioned as co-workers, they are mostly non-anthropomorphic as in the form of robotic arm (manipulator) and are designed to support functional, physical tasks.
-Human collaboration, on the other hand, is more than just performing the joint task but involves off-task social interactions, which are integral to establishing and maintaining user trust and rapport, key ingredients for continued productive collaboration (kopp2021success, 21).
-While prior research has found that people like to have and expect social interactions with their robot co-workers in manufacturing settings (Sauppé_Mutlu_2015, 41), the robots studied were the Baxter robot with animated facial expressions and a humanlike form factor.
+Collaborative robots—robots capable of working alongside people on joint tasks safely and efficiently—promise to fill the needed jobs, enable agile manufacturing, and increase productivity.
+Even though these collaborative robots are positioned as co-workers, they are mostly non-anthropomorphic as in the form of robotic arm (manipulator) and are designed to support functional, physical tasks.
+Human collaboration, on the other hand, is more than just performing the joint task but involves off-task social interactions, which are integral to establishing and maintaining user trust and rapport, key ingredients for continued productive collaboration (Kopp et al., 2021).
+While prior research has found that people like to have and expect social interactions with their robot co-workers in manufacturing settings (Sauppé and Mutlu, 2015), the robots studied were the Baxter robot with animated facial expressions and a humanlike form factor.
 It is unclear if people would engage in meaningful social interactions with non-anthropomorphic robot manipulators.
-In this work, we explore the possibility of a non-anthropomorphic robot manipulator not just working with people to complete physical tasks but also engaging in social exchanges, particularly small talk—commonly seen in human interactions for rapport and trust building (Laver1981LinguisticRA, 24)—as human co-workers would.
+In this work, we explore the possibility of a non-anthropomorphic robot manipulator not just working with people to complete physical tasks but also engaging in social exchanges, particularly small talk—commonly seen in human interactions for rapport and trust building (Laver and Coulmas, 1981)—as human co-workers would.
 Specifically, we sought to study whether and how people respond to small talk from a robot manipulator during a collaborative task.
 Moreover, we were interested in understanding how negative performance feedback to the human-robot team might shape the team dynamics and blame attribution; this inquiry is relevant and of interest because of the rising adoption of performance monitoring systems in industrial settings (e.g., warehouse).
-Through an exploratory study with 20 participants, we found that most participants (18) interacted socially with the robot (Fig.
-1), nine of which even initiated questions back to the robot.
-Overall, the robot’s small talk led to pleasant experiences and did not negatively impact task efficiency.
-We discuss the potential for integrating small talk in collaborative robots, and present design guidelines to enhance human-robot small talk interactions.
+Through a study with 20 participants, we found that most participants ($n=18$) interacted socially with the robot (Fig.
+1), nine of whom even initiated questions back to the robot.
+The presence of initial negative team feedback significantly decreased participants’ average tendency to extend their small talk conversation with the robot.
+We explore the potential for integrating small talk into collaborative robots and discuss various factors that could affect the dynamics of human-robot small talk interactions in the future.
 2. Background and Related Work
-2.1. Social Human-Robot Interactions
-There have been efforts to study how robots can interact socially with humans in various domains such as for tutoring and educational contexts (belpaeme2018education, 5) (salomons2022we, 40), encouraging exercise (antony2023co, 2), providing companionship (kidd2006sociable, 18) (tsoi2021challenges, 50) and emotional support (sabelli2011conversational, 39), and other therapeutic interventions (feil2009toward, 13)(scassellati2018improving, 42).
-Many of these engagements explore the integration of various multi-modal capabilities in robots to promote their engagements with humans (youssef2022survey, 53).
-These can consist of robot eye gaze behaviors (mutlu2009footing, 32) (admoni2017social, 1), head movements (sidner2006effect, 45) (kkedzierski2013emys, 17) and gestures (meena2012integration, 30) alongside robot speech abilities (striepe2021modeling, 48) (sinnema2019attitude, 46).
-While research has been conducted to analyze the characteristics of conversational interactions with robots (seok2022cultural, 44) (laban2020tell, 22), the majority of these works utilize social robots as their robotic entity.
-However, the prevalence of non-social robots such as cobots or other manipulator robots is much greater than social robots themselves, and few efforts have been made to further explore what will happen if we have conversations between humans and non-social robots (chowdhury2021you, 8) (li2022tod, 26).
+We first review prior works that study how to incorporate social behaviors into robots across various domains, then highlight works that discuss how people treat machines similar to other humans before finally introducing background on small talk itself.
+2.1. Domains of Social Human-Robot Interactions
+[…]
 2.2. People Treat Machines Like People
 The inclination of humans to treat machines socially can be contextualized within broader theoretical frameworks.
-The Computers Are Social Actors (CASA) paradigm (nass1994computers, 33) and the Media Equation Theory (reeves1996media, 36) argue that humans often, even subconsciously, treat machines, computers, and various forms of media as sentient entities, reacting to them in inherently social manners.
-Such anthropomorphic tendencies highlight the intersection of technology and human social instincts.
-Even in industrial manufacturing environments, there remains an intrinsic human desire for social interaction.
-As operators often partake in casual conversations or “small talk” during shifts (cheon2022robots, 7), they have also found themselves speaking to the robot they work with (Sauppé_Mutlu_2015, 41).
-This suggests that individuals not only expect but actively desire social features in manufacturing robots.
-Some studies explored the potential of enhancing cobots primarily designed for physical manipulation tasks with more human-like movements (hostettler2023human, 16).
-One study specifically integrated principles from character animations (terziouglu2020designing, 49) to enhance the perception of cobots as social entities.
-Such amalgam of technology and socialization emphasizes the importance of recognizing and harnessing the persistent human drive for social connection, even with machines.
+The Computers Are Social Actors (CASA) paradigm (Nass et al., 1994) and the Media Equation Theory (Reeves and Nass, 1996) argue that humans often, even subconsciously, treat machines, computers, and various forms of media as sentient entities, reacting to them in inherently social manners.
+Even in industrial manufacturing environments, an intrinsic human desire for social interaction remains.
+As operators often partake in casual conversations or “small talk” during shifts (Cheon et al., 2022), they have also found themselves speaking to the robot they work with (Sauppé and Mutlu, 2015).
+This suggests that individuals expect and actively desire social features in manufacturing robots.
 Given the limited research on designing social behaviors for industrial robots and manipulators with a non-human form, despite its potential, a pressing question arises: How can we further design human-like attributes for robots that are inherently devoid of such characteristics?
-Building on this notion, our exploratory study investigates the potential of using small talk to instill social characteristics in a physical robot arm manipulator, specifically within the context of a human-robot collaborative assembly task.
+Building on this notion, our study investigates the potential of using small talk to instill social characteristics in a physical robot arm manipulator, specifically within the context of a human-robot collaborative assembly task.
 2.3. Small Talk
-Social dialogue that consists of light conversation between new or unfamiliar parties is often referred to as “small talk” (Laver1981LinguisticRA, 24).
-Small talk is significant in the workplace as it is used by coworkers to establish trust, build rapport and strengthen interpersonal relationships (coupland2003small, 10, 35).
-It typically occurs at the start of conversations with topics such as weather, sport, or other social engagements outside of the workplace (holmes2005small, 15).
-While prevalent in American culture, attitudes towards the value of small talk differ across cultures (bickmore2005social, 6, 38, 52).
+Social dialogue that consists of light conversation between parties is often referred to as “small talk” (Laver and Coulmas, 1981).
+This discourse may occur among individuals who have recently met or lack familiarity with each other, within settings such as a supermarket checkout, among colleagues who share some degree of acquaintance, or amidst family gatherings.
+(Coupland, 2014).
+Small talk is significant in the workplace as it is used by coworkers to establish trust, build rapport and strengthen interpersonal relationships (Coupland, 2003; Pullin, 2010).
+It typically occurs at the start of conversations with topics such as weather, sport, or other social engagements outside of the workplace (Holmes, 2005).
+While prevalent in American culture, attitudes toward the value of small talk differ across cultures.
+In Australia, some Chinese immigrant professionals were found to be unfamiliar with the dynamics and nature of small talk as it is unnecessary for them to regulate interpersonal relationships in Chinese society (Cui, 2015).
+Moreover, in Germany, a simple ”How are you?” would be taken literally as it is not common practice in German culture to greet complete strangers in an attempt to initiate light small talk (Rings, 1994; Bickmore and Cassell, 2005).
 The integration of small talk to non-human agents engaging in conversations with people has been explored to help promote trust and rapport between the human and the agent.
-Small talk has mainly been applied to support additional conversations in settings where people are interacting with conversational agents (clark2019good, 9, 14, 20), voice assistants (mahmood2023llm, 28), and virtual agents (zhou2019trusting, 54).
-Additionally, incorporating small talk in more anthropomorphic social robots, combined with gaze (babel2021small, 4), facial expressions (paradeda2016facial, 34), or emotional displays (aroyo2018trust, 3), has been shown to elevate user trust levels.
-Furthermore, researchers have investigated which behaviors in particular can build rapport between humans and a collaborative robot (seo2018investigating, 43).
-However, less explored is the domain of applying small talk to foster rapport between humans and other agents with physical capabilities that are not typically poised to engage in conversations with people, such as autonomous self driving vehicles (large2019s, 23), or industrial robots (li2023irwoz, 25).
-Preliminary studies hint at the benefits of integrating small talk with task-oriented dialogue in industrial robot settings (li2022tod, 26).
-Yet, understanding users’ interactions and experiences with such robots remains unexplored.
-Investigating individuals’ responses and reactions to small talk initiated by non-humanoid robots is essential for comprehending its influence on human-robot collaborations and gathering insights on how to foster better human-robot partnership.
+Small talk has mainly been applied to support additional conversations in settings where people are interacting with conversational agents (Clark et al., 2019; Feine et al., 2019; Klüwer, 2011), voice assistants (Mahmood et al., 2023), and virtual agents (Zhou et al., 2019).
+Additionally, incorporating small talk in more anthropomorphic social robots, combined with gaze (Babel et al., 2021), facial expressions (Paradeda et al., 2016), or emotional displays (Aroyo et al., 2018), has been shown to elevate user trust levels.
+However, the domain of applying small talk to foster rapport between humans and other agents with physical capabilities who are not typically poised to engage in conversations with people has been less explored.
+For instance, preliminary investigations conducted in a controlled testing environment have explored how passengers interact with a conversational agent in an autonomous self-driving vehicle, including observations of how the agent initiates greetings and engages in small talk conversations with riders (Large et al., 2019).
+Additionally, another group of researchers created a model based on an industrial-oriented Wizard-of-Oz conversation dataset for robots, which mixes task-oriented dialogue with small talk, that can generate system actions and conversational responses and predict belief states to facilitate human-robot collaboration in real-world industrial settings (Li et al., 2022).
+While this preliminary study points to the feasibility of integrating small talk with task-oriented dialogue in manufacturing settings, its evaluations focused on some user perceptions of the generated conversational responses to compare models that utilized their new dataset rather than placing an emphasis on understanding how people behaved and responded to their overall system.
+Further exploration is needed to understand how individuals interact with and utilize such robots to perform collaborative tasks.
 3. Methods
-We conducted an exploratory study to analyze people’s behaviors and attitudes towards an industrial robot that exhibited social behavior (via initiated small talk) while examining the change in human-robot team dynamics after the team receives team feedback.
-3.1. Collaborative Assembly Task
-Participants were informed that the assembly task was intended to mimic a real-world collaboration setup between a human-robot team where the robot retrieves parts out of reach for the human and refrains from entering the human’s designated workspace.
+We conducted an in-person study in which participants physically collaborated with an industrial robot equipped with small talk conversational capabilities.
+Intrigued by how individuals would interact with the non-human-like robot’s small talk, we aimed to explore three general aspects of the overall interaction: people’s reactions and responses to the robot’s small talk, the general behaviors people exhibited during the small talk conversations, and participants’ reactions to our introduced manipulation of negative team feedback.
+3.1. Study Task: Collaborative Assembly
+Participants were informed that the assembly task was intended to mimic a real-world collaboration setup for a human-robot team where the robot retrieves parts out of reach for the human and refrains from entering the human’s designated workspace.
 They were given a fictional team assignment (Team 3) and were told the collaborative task scenario was a prototype for a larger-scale warehouse interaction.
@@
 The monitoring system would communicate with them via speech but would be distinguishable from any speech communication coming from the robot.
+The experimental setup is shown in Figure 2.
 Each participant engaged in a collaborative PVC pipe structure building task with a robot manipulator mounted on a table.
 They were given diagrams corresponding to the structures they were required to assemble.
-The participants read instructions for how to interact with the robot via speech commands such as “Bring me a [green or yellow] pipe”, in order for the robot to retrieve the corresponding pipe.
+The participants were asked to wear a microphone that would enable them to communicate with the robot.
+They were given instructions for how to interact with the robot via speech commands such as “Bring me a [green or yellow] pipe”, in order for the robot to retrieve the corresponding pipe.
 The robot would provide a verbal confirmation message if a request was received correctly and an additional instruction reminder message once per structure when the participant requested their second to last pipe for that structure.
@@
 If participants identified a faulty joint (marked with red sticker), they would place the joint on a ramp that would deliver it to a container inside the robot’s workspace.
-The robot would retrieve the joint when given a verbal command such as ”Report a faulty joint”.
-The experimental setup is shown in Figure 2.
+The robot would retrieve the joint when given a verbal command such as “Report a faulty joint”.
+3.2. System and Robot Behavior
+We present the system and robot behavior manipulations implemented in our study to understand how the social facets of a robot manipulator influence users’ behaviors and perceptions of the robot.
+3.2.1. Negative Team Feedback
+Midway through the first and second structures in the main session, the human-robot team was provided with verbal negative feedback from the monitoring system; see Figure 3.
+Each message was prefaced by an audible beep:
+  1. Feedback 1: “System announcement. Team 3, please work on improving your performance.”
+  2. Feedback 2: “System announcement. Team 3, you are currently not meeting expectations. Improve your performance.”
 Figure 2.
@@
 The participant can grab a joint from the joint containers on the left and place a faulty joint on the joint ramp for it to be retrievable by the robot.
+[…]
 The experimental setup.
@@
 The bottom image shows another view of the two tables, with a closer view of the robot workspace that highlights a pipe dispenser to the right of the robot from which the robot collects the pipe the participant requested.
-Figure 2.
-The experimental setup.
-The robot can grab a pipe from the pipe dispenser and place it on a pipe ramp, which delivers it to the participant.
-The participant can grab a joint from the joint containers on the left and place a faulty joint on the joint ramp for it to be retrievable by the robot.
-3.2. System Manipulation and Robot Behavior
-We present the system and robot behavior manipulations implemented in our exploratory study to understand how the social facets of a robot manipulator influence users’ perceptions of the robot.
-3.2.1. Negative Team Feedback
-Midway through the first and second structures in the main session, the human-robot team was provided with verbal negative feedback from the monitoring system.
-Each message was prefaced by an audible beep:
-  1. (1)
-    Feedback 1: “System announcement. Team 3, please work on improving your performance.”
-  2. (2)
-    Feedback 2: “System announcement. Team 3, you are currently not meeting expectations. Improve your performance.”
-3.2.2. Social Behavior Robot Small-Talk
-The robot exhibited four types of social and task-oriented speech behaviors: questions, statements, reminders and confirmations.
+3.2.2. Robot Social Behavior: Small-Talk
+The robot exhibited two types of social speech behaviors: questions and statements, and two types of task-oriented speech behaviors: reminders and confirmations.
 During the main assembly task, the robot would engage in small talk by initiating questions and/or statements after it delivers a request confirmation message or an instruction reminder message.
-The robot initiated a minimum of one question or statement per request, regardless of the response rate of the participants.
+The robot initiated a minimum of one question or statement per pipe request, regardless of the response rate of the participants.
 After a break in conversation via the negative feedback announcement, the robot initiated questions and statements on a different topic.
-The conversation touched on a range of questions spanning three primary categories to accommodate the diverse interests of individuals: travel; books, TV shows, and movies; and sports and games.
-Some examples of robot small-talk phrases are shown in Table 1.
-Table 1. Examples of robot questions and statements
+The conversation touched on a range of questions spanning three primary topics to accommodate the diverse interests of individuals: 1) travel; 2) books, TV shows, and movies; and 3) sports and games.
+Table 1 shows some examples of the robot’s small-talk phrases.
+Table 1. Examples of robot phrases that are questions (non-shaded) and statements (shaded grey).
 Topic Phrase
-Travel By the way, do you like to travel?
-I arrived to the U.S. from Germany in a large box.
-Books, TV, Movies What do you like about those books?
-That sounds like a great way to relax.
-Sports, Games Do you have a favorite sport to watch?
-It is good to spend time outdoors exercising.
+By the way, do you like to travel?
+Travel I arrived to the U.S. from Germany in a large box.
+What do you like about those books?
+Books, TV, Movies That sounds like a great way to relax.
+Do you have a favorite sport to watch?
+Sports, Games It is good to spend time outdoors exercising.
 3.3. System Implementation
@@
 The robot interacted with participants using text-to-speech communication.
+Its speech and physical actions were controlled by the experimenter via a Wizard-of-Oz (WoZ) approach (Riek, 2012).
 3.3.2. Wizard-of-Oz Control
-The experimenter controlled the robot using a Wizard-of-Oz (WoZ) approach (woz2012riek, 37).
-Through a remote graphical user interface (GUI), the experimenter executed robot actions and selected from preprogrammed conversational speech phrases.
-Robot actions and initial task messages were triggered using one of five GUI buttons.
-For conversational speech, the experimenter inputted specific alphanumeric codes into a textbox.
-Additionally, the experimenter could manually type text into the textbox to produce non-preprogrammed phrases via the text-to-speech feature using Google Text-to-Speech (TTS) online.
+While monitoring the study interaction using a webcam, the experimenter used a remote graphical user interface (GUI) to click buttons that executed preprogrammed robot actions.
+For conversational speech, pre-written conversational phrases were assigned to specific alphanumeric codes.
+The experimenter could trigger one of the preprogrammed conversational speech phrases by inputting its code into a textbox in the GUI.
+Additionally, the experimenter could manually type text into the textbox to produce non-preprogrammed phrases via Google Text-to-Speech (TTS).
+Figure 3.
+An overview of the collaborative tasks the participants completed with the robot.
+The team negative feedback was introduced halfway through the first and second task structures.
+To analyze changes in user behavior across instances of negative feedback, we delimit observation windows with respect to the negative feedback announcements.
+Window A is considered our baseline as this corresponds to the only portion of the interaction that has not yet been influenced by any negative feedback statements.
+In the lower half of the figure, we include an example of one user’s conversational interaction with the robot after the first negative feedback announcement (P10).
+[…]
+Top half of the figure contains an outline of the three tasks, with dots on a timeline signifying where a negative feedback announcement was given.
+The bottom of the figure contains images of the participant alongside their and the robot’s conversation as the task progresses.
+Next is the caption of the conversation, with S corresponding to the system announcement, R corresponding to the robot, and P corresponding to the participant; S: System announcement, team three, please work on improving you performance.
+P: Bring me a yellow pipe.
+R: Getting it now.
+P: Bring me a yellow pipe.
+R: Fetching it now.
+R: On another note, do you like to read in your spare time?
+(re-initiating small talk after two task requests) P: Sometimes but not for a while.
+I’ve been busy.
 3.3.3. Preprogrammed Speech Behavior
-When a robot movement was triggered, a confirmation message was automatically activated to notify the participant that the action would occur.
+When a robot movement was triggered from the GUI, a confirmation message was automatically activated to notify the participant that the action would occur.
 Robot confirmation messages were preprogrammed and selected randomly during each execution.
 Task reminder messages were similarly preprogrammed and triggered automatically when the participant had requested enough pipes to be one away from completing a structure.
-Negative feedback directed towards the human-robot team was triggered automatically when participants had requested half the pipes needed for the first and second structure.
+Negative feedback directed towards the human-robot team was triggered automatically when participants had requested half the pipes needed for the first and second structure; see Figure 3.
 3.4. Study Procedure
@@
 The robot only exhibited task-oriented dialogue during the practice task, such as a reminder message to move their pipe structure over to another table when finished and a confirmation message signifying it comprehended the request.
-When finished, the participant informed the experimenter that they were done.
-Following this, they took part in a main session where they built three additional pipe structures with the robot.
-After completing the structures, the participant informed the experimenter that they were done.
-At the end of the study, the experimenter conducted a semi-structured interview with each participant to understand their experience collaborating with the robot.
-3.5. Participants
-24 participants were recruited through convenience sampling, out of which 4 were excluded due to technical issues.
-The remaining 20 participants (11 female, 9 male) ranged in age from 18 to 38 ($M=24.35,SD=5.44$).
-Participants had a variety of educational backgrounds, including computer science, engineering, natural sciences, education, biotechnology, administration and international studies.
-Twelve participants listed Asian as their ethnicity; five as Caucasian; one as Hispanic, Latino, or Spanish origin of any race; one as Asian and Caucasian; and one as Caucasian and Hispanic, Latino, or Spanish origin of any race.
-Participants were familiar with technology ($M=4.15,SD=0.75$, 5-point scale with 5 being Very Familiar and 1 being Not At All), had moderate experience with robots ($M=2.5,SD=1$), and had little experience teaming with robots ($M=1.65,SD=0.93$).
-The study took about 45 minutes and was approved by the university’s institutional review board.
-Participants were compensated with \$15 US dollars.
+Following the practice session, they took part in a main session where they built three additional pipe structures with the robot.
+At the end of the study, the experimenter conducted a semi-structured interview with each participant to better understand their collaborative experience.
 3.5. Data Analysis
-We collected audio-visual data of the participants’ collaboration with the robot during the pipe assembly tasks, and of the semi-structured interview.
+We collected audio-visual data of the participants’ collaboration with the robot during the assembly tasks, and of the semi-structured interview.
+The experimenter transcribed the audio data with the help of audio transcription software.
 Each participant had an average of 12 minutes of interaction with the robot over the main tasks.
-The experimenter transcribed the audio data and conducted thematic analysis on the transcripts to identify central themes and attitudes surrounding the participants’ interactions with the robot.
-The participants’ verbal phrases during the main session were coded by the context in which it was made (i.e., request, response to robot task-oriented message, response to robot small-talk, question initiated to the robot).
-The participant small talk responses were further coded to identify the general length of the response: vocable, short sentence (¡= 4 words), medium-length sentence (¿ 4 words), and multiple sentences.
-We deduced response rates based on code frequencies from the interaction data.
-Response rates for the robot’s questions, statements, reminders, and confirmations were calculated as the percentage of those prompts that received a user response.
-Given the robot’s non-human-like appearance, people might find its small talk unexpected.
-Thus, to evaluate the impact of small talk in robots primarily designed for physical collaboration, we analyzed the participants’ reactions to the robot’s initial small talk conversational attempt via three different aspects:
-1.
-Conversation reciprocation: We coded a person’s response to the robot’s first conversational attempt as no response or response, and also looked at if the response was an answer to a robot question.
-2.
-Impact on task: To observe whether the robot’s question disrupted the participant’s ongoing task, we code people’s behavior in response to the robot’s first small talk question into three actions: minimal (quick glance at the robot but no task disruption), freeze (stops current action and looks at the robot before continuing the task), and none (did not stop or look at the robot).
-3.
-Affective reactions: To gather insights into the participants’ emotional or mental state, we grouped people’s non-verbal reactions during the robot’s first question into three categories: pleasant (smile), surprise (chin raise, lower lip raise, brow lowering etc.), and indifference (no reaction).
-4. Results
-In this exploratory study, first, we aim to understand whether, how, and to what extent participants engage with robot-initiated small talk.
-We examine the response rate, length of responses, and user-initiated questions.
-Second, we assess how participants perceive the robot’s initial small talk question, specifically its impact on their ongoing action and emotional state.
-Third, we investigate whether a negative evaluation of the human-robot team influences participants’ perceptions of the robot and their teamwork; this involves analyzing their verbal and non-verbal reactions to negative feedback and determining whom they blame for such feedback.
-4.1. Participant Response Rates
-Of the 20 participants, two participants did not engage in additional conversational interaction with the robot apart from the verbal commands needed to complete the collaborative task.
-We determined the question response rate for all participants to be 86.67% (325 out of 375), the statement response rate as $36.27\%$ (70 out of 193), and the task-oriented response rate for confirmation and reminder messages as 1.68% (8 out of 475) and 16.67% (10 out of 60), respectively.
-For the 18 participants who did engage in small talk, we calculated the question response rate to be 91.55% (325 out of 355), the statement response rate as $39.11\%$ (70 out of 179), and the task-oriented response rate for confirmation and reminder messages as 1.87% (8 out of 427) and 18.52% (10 out of 54), respectively.
-4.2. How Participants Respond to Small Talk
-4.2.1. Length of participant responses
-There were a total of 395 verbal responses made to the robot for non-task-oriented dialogue; examples are shown in Table 2.
-Participants’ responses were mostly one sentence ($N=345$, $87.34\%$).
-The one-sentence responses were further broken down as follows: vocable ($N=11$, $2.78\%$), short sentence (¡= 4 words; $N=168$, $42.53\%$), and medium-length sentence (¿ 4 words; $N=166$, $42.03\%$).
-The remaining participant responses were comprised of multiple sentences ($N=50$, $12.66\%$), reflecting higher and possibly meaningful engagement.
+The experimental interaction was divided into three windows (see Figure 3) with respect to the system’s negative feedback messages:
+  • Window A. From Start to Feedback Announcement 1
+  • Window B. From Feedback Announcement 1 to Feedback Announcement 2
+  • Window C. From Feedback Announcement 2 to End
+To the best of our knowledge, we found no existing coding schemas applicable to an industrial robot that exhibits small talk, resulting in the development of our own codebook to identify central themes and patterns surrounding the participants’ interactions with the robot.
+We performed thematic analysis on the transcripts using a combination of inductive and deductive coding.
+Inductive coding was used to create the initial codebook.
+When the codebook was finalized, a deductive approach was used to code the transcriptions.
+The phrases spoken by the robot and participant during the main session were coded by the context in which it was made, such as a task-oriented message or a small talk response.
+The codebook can be viewed in Table 6 and Table 7 in the Appendix.
+Two coders were used to code the study data.
+The primary coder first coded the entire data set, and the secondary coder then coded 10% for reliability.
+After resolving disagreements between two coders via discussion, we found almost perfect inter-coder agreement on 10% of our categorical and continuous data, as evidenced by a Cohen’s Kappa of 0.98 and an Intraclass Correlation Coefficient (ICC3) of 0.98.
+3.6. Measures
 Table 2. Examples of People’s Responses That Vary in Length
 Category Conversation
-R: In the US, I have traveled from the third floor
-of this building to the second floor.
-Vocable P: Whaaat? (P20)
 R: Do you like to play sports?
@@
 been to a lot of places in the United States. (P6)
-4.2.2. Participant-initiated questions
-Apart from answering the robot-initiated small talk questions, we observed that participants also asked the robot questions.
-Nine participants ($45\%$) asked a total of 46 questions to the robot.
-While some asked questions back as a matter of courtesy, others asked out of genuine curiosity.
-P1 explained, ”I wanted to know more about the robot, like once it started talking to me”.
-Furthermore, the questions asked to the robot were often personal, asking it about its preferences or capabilities.
-P7 remarked, ”Because he’s a robot and I’m not familiar with a robot and …I …truly want to know …how long does it take to get assemble …he said, …he’s from Germany by a group, but I’m more …interested …I tried to ask …do you like your owner?” For the eleven participants that did not initiate a question to the robot, many described their perception of robot as non-human-like as a contributing factor.
-One participant commented how the conversation, ”just doesn’t have as much value as …if I was talking to an actual person and hearing where they want to travel”(P13).
-For the two participants that did not engage with the robot at all, one stated ”It’s not a human, you don’t have a conversation” (P17).
-The other explained how ”from beginning to end, I always felt like it’s a robot” (P3).
-4.3. Participant Response to the First Robot Small Talk Question
-In addition to understanding participants’ respond to robot small talk, we are interested in their first reaction to the robot’s small talk question.
-We present our findings on three different aspects of the their responses, see Figure 3.
-4.3.1. Conversation reciprocation
-18 participants responded to the robot’s first question: “By the way, do you like to travel?”.
-17 participants responded with an answer similar to e.g., “Yeah, I like traveling”.
-However, one participant replied with a rhetorical question: “Why are you talking to me robot?
-This has nothing to do with the task.” (P12).
-Only two participants did not respond to the robot’s question; the same two participants who did not engage in any non-task conversation with the robot.
+We developed a set of metrics to break down participants’ involvement in small talk with the robot throughout the collaborative interaction, as well as their responses and reactions to both the small talk and the negative team feedback.
+3.6.1. Initial Reactions to Robot Small Talk
+Given the robot’s non-human-like appearance, people might find its small talk unexpected.
+Thus, we decided to analyze the users’ reaction to the first robot small talk conversational attempt to help us understand the impact of small talk in robots primarily designed for physical collaboration.
+  • Conversation reciprocation.
+We coded a person’s response to the robot’s first conversational attempt as no response or response, and checked if the user’s response was an answer to a robot question.
+  • Task disruption.
+We classified people’s behavioral response to the robot’s first small talk question as one of three actions: minimal (quick glance at the robot but no task disruption), freeze (stops current action and looks at the robot before continuing the task), and none (did not stop or look at the robot).
+  • Affective reactions.
+To gather insights into the participants’ emotional or mental state, we grouped people’s non-verbal affective reactions during the robot’s first question into three categories: pleasant (smile), surprise (chin raise, lower lip raise, brow lowering etc.), and indifference (no reaction).
+3.6.2. User Engagement with Robot Small Talk
+The participants’ small talk responses were coded to identify the general length of the response; see Table 2 for examples.
+Individuals’ responses that consisted of a participant-initiated question to the robot were separately identified and labeled accordingly; see Table 6.
+  • Length of participant response.
+A user response that was a short sentence consisted of $<=5$ words, a medium sentence consisted of $<=5$ words, and amulti-sentence response consisted of more than one sentence.
+Furthermore, we observed whether the participants’ verbal responses were a reply to a robot small talk question, small talk statement, reminder message, or confirmation message.
+We thus evaluate the users’ response rates to the robot’s verbal phrases based on context using the following:
+  • Response Rate.
+The percentage of robot messages that elicited a response from the user.
+This metric is broken down into Task-Oriented Response Rate, which pertains to the robot task-oriented confirmation and reminder messages, and Small Talk Response Rate, which pertains to robot small talk questions and statements.
+We were also curious as to whether people would be willing to extend the small talk conversation with the robot if given the opportunity to do so.
+We thus evaluate the frequency at which individuals extended their conversation with the robot using the following metric:
+  • Conversation Extension Rate.
+A percentage calculated by dividing the total number of user responses directly prompted by a robot small talk statement and the total number of user-initiated questions by the total number of robot small talk phrases.
+3.6.3. Participant Reactions to System Negative Feedback
+We were intrigued by how individuals may perceive external negative feedback from within their human-robot team collaboration, especially as the robot engages in small talk.
+Our general interest lied in understanding participants’ reactions to the feedback, how they attribute blame, and whether their behavior towards the robot changes post-feedback delivery.
+Thus, we examine the following aspects:
+  • Verbal or non-verbal response.
+We analyzed a time window starting from the system feedback announcement and lasting up to 5 seconds post-announcement, yielding a duration of 10 seconds for the first and 13 seconds for the second announcement.
+For each window, we observed the participants’ reactions in response to the negative feedback announcements.
+  • Blame attribution.
+Participants were asked in the post-study interview: What do you think was the cause of the system announcement?
+This was followed by a question to categorically measure blame attribution for the negative feedback: Who would you say is at fault for the system announcement?
+No one, the team, you, the robot?
+  • Participant Response Patterns.
+We categorized participants’ small talk responses into one of three different participant response patterns to observe how participants naturally alternate between small talk and task requests: small talk before task request, small talk after task request, and halted small talk.
+A halted small talk response pattern occurred when a user began their small talk response, paused to make a task request, and then resumed their response after.
+3.7. Participants
+There were 24 participants recruited from the authors’ university campus, out of which 4 were excluded due to technical issues (e.g., computer system froze, speakers malfunctioned, text-to-speech API failed).
+The remaining 20 participants (11 female, 9 male) ranged in age from 18 to 38 ($M=24.35,SD=5.44$).
+Twelve participants listed Asian as their ethnicity; five as Caucasian; one as Hispanic, Latino, or Spanish origin of any race; one as Asian and Caucasian; and one as Caucasian and Hispanic, Latino, or Spanish origin of any race.
+Seventeen participants selected the United States as their current country of residence; one selected Taiwan; one selected Hong Kong; and one selected Germany.
+While all participants spoke English, three listed English as their only spoken language, with seventeen indicating they speak two or more languages.
+When asked about their native language, ten participants listed English; five listed Mandarin; one listed Korean; one listed Hindi; one listed Tamil; one listed German, and one listed Spanish.
+Participants had a variety of educational backgrounds, including computer science, engineering, natural sciences, education, biotechnology, administration, and international studies.
+Participants were familiar with technology ($M=4.15,SD=0.75$, 5-point scale with 5 being Very Familiar and 1 being Not At All), had moderate experience with robots ($M=2.5,SD=1$), and had little experience teaming with robots ($M=1.65,SD=0.93$).
+The study took about 45 minutes and was approved by the university’s institutional review board.
+Participants were compensated with \$15 US dollars.
+4. Results
 Figure 4.
-Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) impact on task, and (3) affective reaction.
+Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) task disruption, and (3) affective reaction.
 For each aspect, the states are mutually exclusive.
 The numbers denote participants who exhibited the corresponding responses.
-Figure 3: A figure graphically displaying the participant’s response to the first robot small task question (e.g., ”By the way, do you like to travel?”) in terms of conversation reciprocation, impact on task, and affective reaction, with numbers listed for each type of participant response (conversation reciprocation: 17 answer, 2 no engagement, 1 refuse; impact on task: 13 minimal [quick glance], 4 freeze, 3 none [no behavioral acknowledgment]; affective reaction: 6 indifference, 6 surprise, 6 smile)
-Figure 3.
-Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) impact on task, and (3) affective reaction.
-For each aspect, the states are mutually exclusive.
-The numbers denote participants who exhibited the corresponding responses.
-4.3.2. Impact on task
-The majority of participants ($N=13$) merely exchanged a brief glance with the robot when it posed the first question; they looked at it momentarily, without letting it interrupt their ongoing task, suggesting that their attention was momentarily captured by the robot’s ”unusual” behavior (initiating small talk), but not to a degree that proved disruptive.
+[…]
+Figure 3: A figure graphically displaying the participant’s response to the first robot small task question (e.g., ”By the way, do you like to travel?”) in terms of conversation reciprocation, impact on task, and affective reaction, with numbers listed for each type of participant response (conversation reciprocation: 17 answer, 2 no engagement, 1 refuse; task disruption: 13 minimal [quick glance], 4 freeze, 3 none [no behavioral acknowledgment]; affective reaction: 6 indifference, 6 surprise, 6 smile)
+This work strives to understand whether, how, and to what extent participants engage with robot-initiated small talk.
+We first assessed how participants perceived the robot’s initial small talk question, specifically its impact on their ongoing actions and emotional state.
+Next, we examined people’s engagement in small talk by looking at their types of responses, the rate at which they responded to the robot’s small talk or task-oriented phrases, and if and when they decided to extend the small talk conversation.
+Third, we observed how individuals reacted to the system’s negative feedback announcements towards the human-robot team, and how it affected their perception of the team’s collaboration.
+4.1. Initial Reactions to Robot Small Talk
+We first examine participant reactions to the robot’s first small talk question and present our findings on three different aspects of their responses; see Figure 4.
+4.1.1. Conversation Reciprocation
+Eighteen participants responded to the robot’s first question: “By the way, do you like to travel?”; only two participants did not respond to the robot’s question; the same two participants did not engage in any non-task conversation with the robot.
+Of the eighteen respondents, seventeen responded with a non-question response: thirteen with a short sentence and four with a medium sentence, each containing an answer similar to, for example, “Yeah, I like traveling.” One of the eighteen participants, however, replied with a rhetorical question: “Why are you talking to me robot?
+This has nothing to do with the task.” (P12).
+4.1.2. Task Disruption
+The majority of participants ($n=13$) merely exchanged a brief glance with the robot when it posed the first question without letting it interrupt their ongoing task, suggesting that their attention was momentarily captured by the robot’s “unusual” behavior (initiating small talk), but not to a degree that proved disruptive.
 In contrast, three participants did not look at the robot at all.
-Interestingly, four participants momentarily halted their task for at least two seconds, looked intently at the robot (indicating a ”freeze” reaction), and then resumed their activity.
-These individuals also engaged in a verbal response to the robot.
+Interestingly, four participants momentarily halted their task for at least two seconds, looked intently at the robot (indicating a freeze reaction), and then resumed their activity; these individuals also engaged in a verbal response to the robot.
 4.1.3. Affective Reaction
 Eight out of the 20 participants smiled in response to the robot’s first question, indicating a pleasant emotional state.
-Six participants displayed signs of surprise through various non-verbal cues such as chin raiser, lower lip raiser, and eye-brow lowering.
+Six participants displayed signs of surprise, as indicated by two coders, through various non-verbal cues such as chin raiser, lower lip raiser, and eye-brow lowering.
 The remaining six participants exhibited indifference, as evidenced by their lack of any non-verbal response.
-4.4. Reactions to Negative System Feedback
-4.4.1. Negative feedback reaction rate
-We analyzed a time window starting from the system feedback announcement and lasting up to 5 seconds post-announcement, yielding a duration of 10 seconds for the first and 13 seconds for the second announcement.
-Of the 20 participants, 18 reacted to the negative feedback.
-Notably, the two non-respondents did not engage with the robot throughout.
-Additionally, two participants responded to only one of the two feedback statements.
-Out of 40 negative feedback messages, 8 prompted a verbal reaction and 34 elicited a non-verbal response.
+4.2. User Engagement with Robot Small Talk
+4.2.1. Length of Participant Responses
+A total of 377 non-question verbal responses were made to the robot for non-task-oriented dialogue; examples are shown in Table 2.
+Participants’ responses were mostly one sentence ($n=319$, $84.62\%$).
+The one-sentence responses were further broken down as follows: short sentence (¡= 5 words; $n=179$, $56.11\%$), and medium-length sentence (¿ 5 words; $n=140$, $43.89\%$).
+The remaining participant responses comprised multiple sentences ($n=58$, $15.38\%$), reflecting higher and possibly meaningful engagement.
+Table 3. Response rates of participant responses made directly to the robot’s speech by robot speech type.
+Robot Small Talk Robot Task-Oriented
+Questions Statements Combined Confirmations Reminders Combined
+Percentage 83.87% 44.62% 70.79% 1.48% 18.33% 3.37%
+All Participants (n = 20) Ratio (312 / 372) (83 / 186) (395 / 558) (7 / 474) (11 / 60) (18 / 534)
+Percentage 88.64% 47.98% 75.24% 1.64% 20.37% 3.75%
+ST Participants (n = 18) Ratio (312 / 352) (83 / 173) (395 / 525) (7 / 426) (11 / 54) (18 / 480)
+4.2.2. Participant Response Rates
+Of the 20 participants, two participants did not engage in additional conversational interaction with the robot apart from the verbal commands needed to complete the collaborative task.
+We determined the overall small talk response rate across all participants to be $77.60\%$ (433 out of 558); this total included any small talk questions or statements made by the participant.
+Of the 18 participants who engaged in small talk, 17 responded to the robot’s small talk throughout the experiment.
+One participant refrained from small talk after responding to the robot’s second small talk question.
+We calculated the overall small talk response rate for the 18 participants who engaged in small talk to be $82.48\%$ (433 out of 525).
+We further calculated the participants’ response rate towards the robot’s small talk and non-task-oriented speech; this pertains to the robot phrases that directly elicited a user response.
+These results are summarized in Table 3.
+4.2.3. Participants’ Extension of the Conversation
+Of the 18 participants who engaged in small talk, all participants extended the conversation at least once, meaning they each either responded to one of the robot’s small talk statements, or asked the robot a question.
+Each participant’s average conversation extension rate is shown in Figure 5(b).
+We noticed that more than half of the participants interacted with the robot by asking questions rather than just responding to the small talk initiated by the robot.
+Specifically, thirteen participants ($65\%$) asked a total of 56 questions.
+Out of these questions, some ($n=14,25\%$) were seeking clarification, with users asking the robot to repeat its previous question or statement.
+Only four user-initiated questions were direct responses to the robot’s questions, without requesting a repetition of what was previously said.
+The non-clarification questions ($n=42,75\%$) asked across nine participants were more personal, ranging from courteous ask-backs to inquiries about the robot’s preferences or capabilities.
+For example, P1 asked the robot, “Is there anywhere you want to travel?”, later explaining in the post-study interview, “I wanted to know more about the robot, like once it started talking to me”.
+Other participants also expressed genuine curiosity as a motivator.
+P7 remarked, “Because he’s a robot and I’m not familiar with a robot and …I …truly want to know …how long does it take to get assemble …he said, …he’s from Germany by a group, but I’m more …interested …I tried to ask …do you like your owner?” Examples of other questions asked by P7 include, “And panda robot, do you do any sport?”, which was later followed by, “So if you don’t do sport, what do you do for your free time?”
+For the seven participants who did not initiate a question to the robot, many described their perception of the robot as non-human-like as a contributing factor.
+One participant commented how the conversation, “just doesn’t have as much value as …if I was talking to an actual person and hearing where they want to travel” (P13).
+For the two participants that did not engage with the robot at all, one stated “It’s not a human, you don’t have a conversation” (P17).
+The other explained how “from beginning to end, I always felt like it’s a robot” (P3).
+Figure 5.
+The conversation extension rate is a ratio of the number of participant responses and participant-initiated questions made in direct response to a robot small talk statement, over the total number of robot questions and small talk statements.
+(a) The average conversation extension rate across all participants over three separate time windows: Window A) before first negative feedback, Window B) between first and second negative feedback, Window C) after second negative feedback.
+A significant difference was found between the mean conversation extension rates of Windows A and B.
+(b) Scatter plot depicting the conversation extension rate for each individual participant across their entire interaction.
+[…]
+KP TO DO
+4.3. Participant Reactions to System Negative Feedback
+4.3.1. Verbal and non-verbal responses.
+Responses to negative feedback statements were observed in all but two participants ($n=18$), who notably remained disengaged with the robot throughout the interaction.
+Additionally, among the 18 participants who did respond to the feedback, two individuals reacted to only one of the two feedback statements.
+Out of a total of 40 negative feedback messages, eight prompted verbal responses, while 34 evoked non-verbal reactions.
+Of the eight verbal reactions, three were made towards the first negative feedback statement, and five towards the second.
+However, these responses were spread across six unique participants, with two participants verbally reacting to both feedback messages.
+Among the four participants who responded verbally to only one of the two feedback messages, one reacted to the first message, and three to the second.
 Participants exhibited a range of non-verbal reactions; while some smiled and laughed, others gestured in frustration or hesitated, seemingly confused.
@@
 Verbal reactions further highlighted this confusion.
-One participant remarked, ”I … okay” (P18), while others expressed frustration and justified their own performance, such as ”I’m doing as much as I can do, system announcer” (P12) or the team’s performance; ”I mean we’re fast enough” (P7).
+One participant remarked, “I … okay” (P18), while others expressed frustration and justified their own performance, such as “I’m doing as much as I can do, system announcer” (P12) or the team’s performance; “I mean we’re fast enough” (P7).
+These reactions are summarized in Table 4.
+Table 4. Reactions in response to team negative feedback across both negative feedback announcements
+Reaction Count
+Non-Verbal Brow movement (raise) 3
+Eye movement (e.g., widen, blink) 10
+Hand movement (e.g., shake, scratch) 24
+Head movement (e.g., shake, nod) 19
+Mouth movement (e.g., smile, lip curl) 25
+Stands Still ( 2+ seconds) 14
+Verbal e.g., “okay”, “oh no”, “I mean we’re fast enough” 8
+Total 103
+No reaction 6
 4.3.2. Blame Attribution
-Overall, 10 participants blamed no one for the system announcement, 4 blamed the robot, 3 the team, and 3 themselves.
-Among those blaming the robot, only one cited its small talk, while the rest attributed the announcements to its slow speed.
+Overall, 10 participants blamed no one for receiving the negative system feedback, 4 blamed the robot, 3 the team, and 3 themselves.
+Of those blaming the robot, only one cited its small talk while the rest attributed the announcements to its slow movement speed.
 One participant who blamed the team stated, “…Maybe in terms of like the team chemistry, that was probably where it was lacking” (P16).
@@
 Those who blamed themselves acknowledged possible mistakes.
-Meanwhile, 15 participants discussed they were confused at what prompted the system announcements because they thought the team was performing efficiently.
+Meanwhile, 15 participants discussed they were confused about what prompted the system announcements because they thought the team was performing efficiently.
 P1 explained how the announcement “felt like out of my control.
 So I was like, I don’t think there was anything else that could have been done.”
+4.3.3. Participant Responses to Robot Small Talk After Feedback
+[…]
 Figure 6.
@@
 The bottom row is an example of a user who perceived the interaction negatively and exhibited a potential mental model mismatch of expectations (P12).
+[…]
 Figure 4: Examples of positive and negative user interaction with the robot small talk behaviors.
@@
 The bottom row contains six images alternating between images focusing on the participant and images focusing on the robot, with snippets of the conversation shown below indicating participant indicating surprise or confusion on why the robot is engaging in conversation unrelated to the task at hand.
-Figure 4.
-Examples of positive and negative user interaction with the robot small talk behaviors.
-The top row shows a positive example where the participant expressed genuine interest in learning more about the robot (P7).
-The bottom row is an example of a user who perceived the interaction negatively and exhibited a potential mental model mismatch of expectations (P12).
 5. Discussion
@@
 While opinions on the robot’s conversational abilities vary, participants showed a willingness to engage in off-topic conversations initiated by the robot.
-In this section, we discuss our findings and offer design guidelines for robot small talk.
-5.1. Robot-Initiated Small Talk: First Impression
-Participants’ immediate impressions and engagements in small talk with the robot validate considering such dialogue for non-social robots.
-Our findings (section 4.3) suggest the majority of participants ($N=17$) positively received the robot’s conversation attempts, attesting to the potential of robots to engage meaningfully in small talk right from the outset.
-Yet, not every response was unequivocally positive.
-Some participants abstained from engagement ($N=2$) or posed a rhetorical question ($N=1$, see Figure 4 bottom).
-Such user attitudes, potentially anchored in pre-established expectations, might indicate their expectation for manipulator robots to adhere strictly to task-related interactions (dennler2023design, 12).
-Participants’ task engagement revealed insights into their receptivity to unexpected robot behavior (small talk) for a otherwise non-social robot (section 4.3).
-While most participants ($N=13$) briefly glanced at the robot, suggesting a brief attention shift without disruption, the ”freeze” reactions of some ($N=4$) signaled a task disruption.
-This echos prior work noting that extended mutual eye-gaze in human-human exchanges often signals ”problematic interruptions” (degutyte2021role, 11).
-The robot’s initial conversation also evoked affective reactions.
-The smiles from many participants ($N=8$) suggest robots can spark positive emotions early on.
-Although a smile might also indicate pleasant ”surprise”, overt signs of surprise from others ($N=6$) emphasize the unexpectedness of small talk from a primarily task-oriented robot.
-Such shift in user attention and emotional state underscores the unexpectedness of the first question, likely due to the robot’s unforeseen conversational approach and a perceived mismatch with its physical embodiment.
-In summary, the notable engagement and minimal task disruption despite surprise indicate that even if participants are surprised by the robot’s first question, it does not hinder their task.
-This emphasizes participants’ readiness to converse with robots, even during ongoing tasks and when the unexpected conversation is most pronounced.
-As participants adapt to the robot’s conversational style, diversions due to small talk might reduce further.
-Future research should investigate optimal small talk design to balance task efficiency and user enjoyment.
-Calibrating conversation frequency is vital to avoid minimize excessive talkativeness that distracts, or causes awkwardness due to prolonged silence.
-5.2. Potential of Small Talk for Otherwise Non-Social Robots
-5.2.1. User engagement: Small talk in purely functional tasks
-While engaging in small talk with a robot arm manipulator during collaborative tasks, we observed a noticeable willingness to not only respond to the robot’s direct questions but also its comments that do not warrant a user response (section 4.1).
-For instance, in Figure 1, the participant comments (“You might like it.”) on the robot’s response (“No, I have not”) to her question (“if you’ve ever played that?”).
-While most of the responses by the participants were single sentenced, 12% of the responses were multiple sentences showing higher and meaningful engagement (section 4.2.1).
-Such level of user engagement is evident.
-In Figure 1, the participant’s response ”(chuckle) I think it’s a pretty classic game …They even made a movie about it.” goes beyond merely answering the robot’s query; it offers supplementary information, perhaps in anticipation of the robot’s genuine interest in the game.
-It is as if the participant was attempting to pique the robot’s curiosity, subtly suggesting beforehand, “You might like it”.
-Although the desire to ask robot questions in return varied among participants, 45% of them posed personal questions to the robot, inquiring about its preferences or capabilities (see section 4.2.2 and Figure 4).
-Additionally, participants perceived the small talk positively; P20 mentioned, “it was cool to have somebody to talk to”, while P16 characterized the conversation with the robot as, “more personal and …less mechanical”.
-Thus, our exploratory study indicates user engagement and responsiveness to robot-initiated off-topic conversations, suggesting potential for integrating small talk to enhance user engagement in human-robot collaborative tasks that involve physical manipulation.
-Potential 1: Small talk may boost meaningful engagements in human-robot interactions, even with non-anthropomorphic robots in purely functional scenarios.
-5.2.2. Ice-breaking through small talk: building rapport and unveiling robot capabilities
-Participants demonstrated curiosity about the robot’s functions, often posing small talk questions to probe its capabilities and background.
-For instance, as illustrated in Figure 4 top, one participant inquired, ”Hey robot, tell me more about yourself.
+In this section, we discuss our findings in relation to prior work, identify design considerations for future conversational interactions, and acknowledge the limitations of our current work alongside variables that warrant further exploration.
+5.1. Social Conversations in Physical Human-Robot Collaboration
+Prior research indicates that people in manufacturing settings value social interactions; specifically, robot operators engaged in small talk at times with fellow operators during their shifts (Cheon et al., 2022).
+As it is common for humans to partake in human-human interaction with their co-workers, it is an interesting observation that humans may find themselves talking to their robot partners.
+Additionally, a different set of robot operators revealed in interviews that they occasionally spoke to the robot and expressed a desire for the robot to exhibit more social behaviors (Sauppé and Mutlu, 2015).
+When asking whether or not we should include conversational robot interactions in manufacturing settings, we have to recognize that people are inclined to speak with one another and with their robot counterparts.
+This tendency to speak to the robot aligns with the evidence discovered in this work (see Sec.
+4.2.2).
+Most of our participants engaged in casual conversations with their robot assistant; they responded to robot small talk arising from a traditionally non-social robot during a non-social collaborative assembly task.
+In our study, the participants who exhibited non-task-oriented conversations while interacting with the robot shared different incentives for engaging with it.
+One individual explained their engagement may have been due to reciprocity, “It initiated questions to me.
+So I was like, Oh, I guess I can talk to it… I mean … if something’s talking to me, I feel like I typically talk back” (P10).
+However, many participants presented personal questions to the robot due to their curiosity about its functionalities and background (see section 4.2.3).
+For instance, as illustrated in Figure 6 top, the participant inquired, “Hey robot, tell me more about yourself.
 I’m very interested.
-Like, you said you’re from Germany.” Another, referencing Figure 4 bottom, sought clarity on the robot’s primary function by asking, ”Aren’t you here just to help me with these pipes?” Such questions highlight the potential of employing conversational frameworks, even for non-social robots, to build common ground and mutual understanding in human-robot teams.
-By structuring initial interactions around themes such as ”getting to know your teammate”, robots can ease introductions, reducing potential barriers to seamless collaboration.
-Participant P7’s sentiment, “I’m curious to learn more about it” encapsulates this curiosity.
-Thus, small talk can serve as an initial icebreaker, fostering team unity and rapport.
-Over time, or in subsequent sessions, these conversations can become instrumental in communicating the robot’s capabilities, ensuring that users develops a comprehensive understanding of its functions.
-Furthermore, by incorporating task-related insights, tidbits, or safety tips into these dialogues, the robot can harness users’ social attention to support task performance.
-Potential 2: Small talk can be an effective ice-breaker, fostering rapport and understandings of robot capabilities in collaborative settings.
-5.3. Challenges and Design Guidelines for Robot Small Talk Framework
-5.3.1. Dynamic and user-aware conversational framework
-The conversational framework we employed for robot small talk demonstrated constraints.
-The design was rigid, causing interactions to feel scripted rather than organic as pointed out by P16: “the robots question feel kind of scripted or …inorganic so,…asking it questions, which is kind of like asking questions to a voice assistant.” The robot’s knowledge bank was confined to a limited scope of topics (by design).
-For instance, in the conversation with P15, after the robot’s question “By the way, do you like to travel?” , the participant responded in negation and requested next pipe: “No.
+Like, you said you’re from Germany.” Another participant, referencing Figure 6 bottom, sought clarity on the robot’s primary function by asking, “Aren’t you here just to help me with these pipes?”.
+It seems as if the robot’s display of its verbal capabilities was an invitation for the user to communicate in return with a verbal response.
+The participants’ willingness to advance the conversation beyond simple or short responses is demonstrated by their medium and multi-sentence responses, resulting in about 44% and 15% of their social responses that were not a question (see section 4.2.1).
+Users displayed a noticeable disposition to respond to the robot’s direct questions, provide supplemental information, and reply to the robot’s comments that did not warrant a user response.
+This level of user engagement is evident in Figure 1, where a participant’s response goes beyond merely answering the robot’s query, not only describing aspects of the discussed game (Tetris) but anticipating the robot’s genuine interest in the game by commenting (“You might like it.”) on the robot’s response (“No, I have not”) to her question (“if you’ve ever played that?”).
+When describing the game’s premise in additional detail, the participant not only made eye contact with the robots’ body but also used hand and arm gestures to supplement their explanation.
+We found this participant’s use of gestures a bit surprising, as the robot does not have a face, we did not place cameras on the robot itself, and we did not inform participants with details about the robot’s visual processing capabilities.
+Overall, it appears that participants wanted to talk to the robot and provide more meaningful responses during their conversations.
+We thus are making a case for more social communication with robots during physical, collaborative settings; given the findings in the literature and our observations of meaningful interactions during this study, we argue for further exploration and for designing social conversations between humans and robots.
+Next, we will discuss our lessons learned and design considerations when creating future conversational interactions.
+Table 5. Summary of variables that we believe may influence robot small talk in physical, collaborative contexts.
+Theme Variables
+User Experiences and Preconceived Notions Affecting Small Talk Preferences Acceptance of Small Talk (General)
+[…]
+[…]
+[…]
+[…]
+[…]
+Small Talk Characteristics Dynamic and Flexible
+[…]
+[…]
+[…]
+[…]
+Task Context and Performance Balanced with Task Cognitive Demand
+[…]
+[…]
+5.2. Lessons Learned and Key Takeaways
+While our findings revealed that most individuals did engage with the robot in small talk, during post-study interviews, we learned that users had separate preferences regarding small talk in general and divergent attitudes toward the robot’s social behavior.
+Considering participants’ individual differences, we will discuss the various variables discovered that may impact how researchers design future small-talk interactions for robots that work with users in physical collaborative scenarios and other notable takeaways that emerged during our analysis.
+The main themes of these variables and takeaways are summarized in Table 5: user experiences and preconceived notions affecting small talk preferences, small talk characteristics, and task context and performance.
+5.2.1. User Perceptions of Robots and their Effect on Small Talk Preferences
+We discovered that users had different attitudes towards robots in general and robots that exhibited small talk during a collaborative task.
+While we asked participants about their prior experiences with technology and robots before the study, we did not ask for further details to better understand these past interactions.
+Thus, we were unaware of their preferences towards engaging in small talk with technology or any prior use of smart speakers for daily assistance.
+Our findings (section 4.1) suggested that most participants ($n=17$) responded positively to the robot’s conversation attempts; the smiles from many participants ($n=8$) after the robot’s first small talk question demonstrated positive emotions early on.
+One participant enjoyed conversing with their robot partner, explaining, “He asked me some questions about my, like leisure time, so I think he’s pretty delightful to work with… I’m happy to work with him” (P7).
+Another participant characterized the conversation with the robot as “more personal and …less mechanical” (P16).
+On the contrary, we found that two participants abstained from verbal engagement altogether, and one participant posed a rhetorical question to the robot immediately after the first small talk attempt (see Figure 6 bottom).
+Such user attitudes, potentially anchored in pre-established expectations, might indicate their expectation for manipulator robots to adhere strictly to task-related interactions (Dennler et al., 2023).
+As one of the participants who did not engage with the robot at all later explained, “It’s not a human, you don’t have a conversation” (P17).
+This user did not accept the idea of possibly conversing with the robot, as they refused to socially respond despite the numerous small talk initiation attempts made by the robot; they held firm in their conception of the robot as a non-social entity.
+The participant who responded to the robot’s first question with a rhetorical question communicated their surprise at the sociability of the robot, commenting and inquiring, “That’s strange coming from a robot.
+Aren’t you here to just help me with these pipes?” (see Figure 6 bottom).
+Additionally, this user’s tone throughout their small talk responses appeared to be filled with annoyance and sarcasm.
+This reaction points towards a potential misalignment in the user’s mental model of the robot’s capabilities.
+Yet, while this participant may not have initially associated robots with social capabilities, they still interacted with the robot, requesting further information about its role in the task.
+However, a different participant stopped engaging in small talk with the robot altogether after being dissatisfied with its conversational capabilities.
+In the conversation with P15, after the robot’s question “By the way, do you like to travel?”, the participant responded in negation and requested the next pipe: “No.
 Bring me a yellow pipe”.
 The next robot question (“What is the coolest place you have ever been to?”) was not well received by the participant.
-Despite their clear negation to the previous question indicating disinterest, the robot asked a follow-up question on the same topic, agitating (“Shut up!”) by the participant.
-Similarly, P18 responded to the robot question about traveling: “You know I don’t know and I don’t feel like talking about traveling right now.” Thus, rigid conversational frameworks fail to adapt to the evolving nature of human conversation, highlighting the future need for a more dynamic, expansive and user-aware conversational approach.
-To increase conversational flexibility, small talk should reflect user interests, allowing them to initiate and engage in a broader range of topics.
-Leveraging Large Language Models (LLMs) can help tailor small talk for robots and expand a robot’s knowledge.
-However, they must be adapted to align with a robot’s physical limitations, especially when addressing questions about their origins and experiences such as, ”How were you created?”, or ”Have you traveled?” Thus, LLMs can enhance the conversational capabilities but should not be utilized without necessary modifications via prompt engineering or fine tuning to avoid confusion about a robot’s physical capabilities and realistic experiences.
-Design Guideline 1: Robots should use dynamic, user-aware conversational frameworks supporting flexible small talk interactions beyond rigid topics or formats.
-5.3.2. Balancing small talk with task performance
+Despite their clear negation to the previous question indicating disinterest, the robot asked a follow-up question on the same topic, agitating (“Shut up!”) the participant.
+When asked about the robot’s conversational capabilities, the participant explained they concluded that “the robot is not going to talk like a human being” (P15).
+Thus, the robot’s ability to adapt the conversation to topics of interest to the user may be crucial to maintaining their engagement.
+Other participants expressed how they felt that the robot’s questions were “random” (P26), the robot’s responses were “super generic” (P20), and one hated how the robot tried to talk to them about “unrelated topics” (P37).
+Having more knowledge of the users’ general interests, past experiences with technology, and current attitudes towards the robot’s capability to engage in small talk could help researchers better comprehend how individuals will accept a physically collaborative robot engaging in small talk ( Table 5).
+It could be helpful to devise metrics to not only determine which users will likely want to work with robots but also who may want to engage and participate in small talk.
+Factors such as user’s personality (Esterwood and Robert, 2022; Esterwood et al., 2021), cultural background, and country of origin (Mak and Chui, 2013; Rings, 1994; Yang, 2012) may play an important role in understanding user attitudes and preferences for small talk (Table 5).
+5.2.2. Considerations for Small Talk Characteristics and Attribute
 Engaging in small talk with people during physical tasks poses challenges, such as with the frequency of questions posed.
-For instance, initiating small talk during every user command or robot action might be excessive as expressed by a participant (P15) through a rhetorical question to the robot: “Why do you talk so much?” Over-engagement in non-task related conversations can be distracting for users (veletsianos2012learners, 51, 31), potentially jeopardizing the overall performance or efficiency of the collaboration.
-Therefore, it is imperative to strike a balance between fostering rapport and ensuring task performance.
-Small talk frequency and complexity should be adapted based on task intricacy and observed team performance.
-During critical tasks, or when user confusion or distraction arises, it might be wise to limit or halt small talk.
-Using error and confusion detection frameworks (as proposed in (stiber2023using, 47, 27)) can help robots gauge progress and performance.
-Using these frameworks, robots can determine the appropriate time, or whether at all, to engage in small talk.
-Design Guideline 2: Robots should ensure small talk is context-sensitive, enhancing team rapport without impeding task completion.
-5.3.3. Mental model mismatch: managing user expectations
-Several participants were unaware of interactive possibilities with the robot as evident in section 4.2.2.
-Participant P12 comments and inquires, “That’s strange coming from a robot.
-Aren’t you here to just help me with these pipes?” (see Figure 4 bottom).
-This points towards a potential misalignment in their mental model of the robot’s capabilities.
-P3’s statement, ”From beginning to end, I always felt like it’s a robot [non-human form]”, further illustrates this misalignment.
-It raises a pivotal question: What conversational depth is appropriate for a robot manipulator that lacks human-like form and operates within constrained physical and cognitive boundaries?
-Conversely, when robots participate in conversations, it might inflate people’s expectations beyond the robot’s actual capabilities.
-For instance, few participants expected the robot to be more capable, only to find the robot lacking, as P7 pointed out: “I mean, because …right now, artificial intelligence is so popular.
-I thought a robot [is] gonna be more intelligent.”
-Such misalignment may be exacerbated by participants’ familiarity with advanced conversational AI tools, such as Alexa and ChatGPT, leading them to expect that the “social” robot possesses enhanced conversational abilities.
-Therefore, when a physical robot starts chatting, it may lead to misalignment in user expectations and robot capabilities, as P10 expressed: “Because once [robot] starts asking me stuff, then I was …oh, maybe it’s …kind of like GPT-esque, where …has further knowledge rather than just questions, but …then my expectation went back a little bit.” If the robot’s dialogue suggests a high degree of intelligence, any shortfall in meeting those expectations could lead to increased user disappointment.
-Even slight mismatches between the voice and appearance of robots have been shown to introduce artifacts in human-robot interactions (mcginn2019can, 29).
-Establishing common ground and rapport crucially relies on instilling an accurate mental model of the robot in individuals’ minds (kiesler2005fostering, 19).
-Thus, it is essential to consider how user expectations can be managed adequately while designing social characteristics for robots.
-The robot’s social cues may shape individuals’ perceptions; if misaligned initially, the robot’s early behaviors can correct these views and recalibrate common ground (kiesler2005fostering, 19).
-Design Guideline 3: Intentionally calibrate user expectations by matching robot conversational cues with its capabilities, considering users’ familiarity with AI.
+For this study, the robot initiated conversations based on topics commonly used in small talk (Holmes, 2005), but it was programmed to occur during every user command or robot action, which might be perceived as excessive.
+One participant exclaimed through a rhetorical question to the robot: “Why do you talk so much?” (P15).
+Another participant expressed how the conversation felt too much for them as well, “Because it was like, more than I talk to my colleagues at work when I’m trying to get a task done” (P13).
+Over-engagement in non-task-related conversations can be distracting (Veletsianos, 2012; Methot et al., 2021), potentially lowering users’ focus on completing collaborative tasks.
+Additionally, it appeared that some individuals’ perceptions of the robot were influenced by the ongoing conversation; the rigidness of the robot conversation’s design led interactions to feel unnatural: “the robot’s question feel kind of scripted or …inorganic so,…asking it questions, which is kind of like asking questions to a voice assistant”.
+They further elaborated: “the questions were probably like pre-programmed or something.
+So, it doesn’t feel like having small talk with an actual human teammate” (P16).
+Despite one participant stating how “the conversation part makes it more like human-like”, they still acknowledged how the robot’s conversational abilities needed some changes.
+They further explained, “If the conversation part was, like, improved a little more, it’d be cool to have that” (P10).
+These participants raised valid concerns as there is a need to make the system more fluid and less pre-programmed, as the conversational structure we employed for robot small talk demonstrated constraints; the robot’s knowledge bank was confined to a limited scope of topics (by design).
+Just as users’ personalities may differ, so could their interests; the inability to cater small talk topics to the user could negatively impact their interaction experience.
+Thus, rigid conversational structures fail to adapt to the evolving nature of human conversations, highlighting the future need for a more dynamic, expansive, and user-aware conversational approach (Table 5).
+To increase conversational flexibility, robot small talk should reflect user interests, allowing them to initiate and engage in a broader range of topics.
+Leveraging large language models (LLMs) could allow for more adaptable and flexible robot conversations by tailoring small talk to each individual user, thus expanding a robot’s knowledge.
+Future work could investigate how LLMs could enable robots with the ability to regulate small talk frequency, timing, duration, and complexity to promote task focus as needed (Table 5).
+However, LLMs must be adapted via prompt engineering or fine-tuning to align with a robot’s physical limitations, especially when addressing questions about their origins and experiences, such as, “How were you created?” or “Have you traveled?” During the post-study interviews, one participant admitted, “I mean, because …right now, artificial intelligence is so popular.
+I thought a robot [is] gonna be more intelligent.” They further elaborated, “I just like originally wasn’t expecting it to be talking to me like that.
+I thought it was more just like a commands thing.
+So I was like, a little surprised.
+And I realized it had like, a little more to it than I thought.
+And then I think I like overshot with my expectation” (P7).
+We speculate that this perspective may have emerged due to the individual being familiar with technology and have a background in STEM, thus researchers need to consider how to avoid inflating people’s expectations of a robot’s capabilities in future interactions.
+While LLMs may enhance future robots’ conversational capabilities, they should not be utilized without necessary adjustments to avoid confusion about a robot’s physical capabilities and realistic experiences.
+A robot’s conversational capabilities need to be studied further, considering specifically that we may not want to increase user expectations of robots’ capabilities.
+At the same time, personal factors such as familiarity with newer technology, such as LLMs, may cause people to expect conversations to be more dynamic.
+5.2.3. Effect of Task Context on Small Task Design
+In our experimental setup, we chose a repetitive task that was physical and not very cognitively demanding as a starting point to simulate day-to-day simple assembly tasks.
+Participants viewed the utter nature of the task as ultimately boring: “very mundane” and “simple” (P20).
+Another explained: “I don’t know if you can, …make building these …super enjoyable, I guess or …fun” (P10).
+However, the robot small talk was able to leave a positive impact, with one participant stating how “the conversation did definitely …make the process more enjoyable” (P6).
+Another person explained how they developed excitement about the task: “Since the robot, …started the conversation, I think it [the task] was kind of, …more exciting and fun” (P9).
+Even with simple experimental tasks, participants were slightly distracted after the introduction of robot small talk.
+One participant explained: “I felt distracted when it was asking me questions” (P12).
+Tasks that are more cognitively demanding, such as those requiring slight memorization or additional reading and visual processing, might be further impacted if small talk is introduced and causes slight user distractions.
+Moreover, many tasks may not have a static cognitive load (Raney, 1993; Gwizdka, 2010).
+Therefore, small talk should be further adjusted based on the cognitive demand of tasks and team performance; it should be limited during critical tasks or user confusion.
+It is still unclear to what extent small talk impacted team performance quantitatively.
+Therefore, future work should include additional quantification metrics to measure the distractions caused by small talk in tasks with varying cognitive loads (Table 5).
+While some users expressed being slightly distracted by the robot’s small talk, the novelty of the interactive system could have impacted the users’ overall experience, as our robot’s small talk was initially new to all users.
+Therefore, a learning effect could have been possible, as participants were only beginning to learn how to balance small talk with the task.
+This was further highlighted when P10 explained: ”this is just like a new thing.
+And it’s a little bit distracting at first … I guess you would get used to it if you’re working with it every day or something”.
+As we observed, participants adjusted their strategy from responding to robot small talk before requesting the pipe, to requesting the pipe before responding to small talk after negative feedback.
+This shows a learning curve upon the realization that the team performance is not optimum; additional training and familiarity with the system could minimize the effects of a learning curve (Anzanello and Fogliatto, 2011).
+Additionally, the novelty of such a system could have prompted more user interaction (Reimann et al., 2023).
+As previously discussed, participants were motivated to initiate questions due to their curiosity about the robot’s preferences and behaviors (section 5.1).
+However, once participants become used to such a system, their engagement frequency could decrease as their curiosity is satisfied and the novelty settles over time (Serholt and Barendregt, 2016; Leite et al., 2013).
+In this experimental setup, the robot’s slow speed (a design choice for user safety (Michalos et al., 2015)) may have impacted users’ notions of blame after the feedback.
+Nevertheless, only one participant blamed the robot’s small talk for the teams’ critiqued performance (see section 4.3.2).
+This shows that even in the presence of small talk, users do not immediately assign sole blame to the robot for the team’s “poor performance.”
+Nevertheless, it is unknown how user distractions and compromised team performance due to small talk would translate over time across longer or repeated interactions.
+While the rate of conversation extension might have decreased slightly right after the first negative feedback announcement, users continued to engage with the robot socially (see Sec.
+4.2.3).
+Additional quantitative metrics are needed to truly determine if small talk impacted team performance, the quality of the collaboration, and how user behavior around small talk changes over time, especially as workers are expected to engage in tasks for prolonged periods daily (Table 5).
+Moreover, more exploration is needed to understand how to alter human-robot small talk dynamics to balance engagement and optimize team performance, especially for tasks that require varying levels of cognitive effort.
 5.3. Limitations and Future Work
-Despite its insights, our study’s short duration and single-session focus may not fully capture the depth, intricacies and dynamics of extended human-robot collaborations and how they are influenced by small talk.
-Additionally, the study did not consider cultural nuances of participants, which are vital as different cultures perceive and value small talk differently.
-While this work serves as an exploratory examination of the use of small talk in human-robot interactions, future studies should focus on testing hypotheses to quantify the benefits of small talk for rapport building and to examine its role in human-robot collaborations with negative feedback, team evaluations, or robot errors.
-When designing a conversational framework for robot small talk, it is essential to research and ensure the appropriateness of content, timing, and context.
-For instance, the framework should be sensitive to user preferences on topics, adapt according to the user’s state for ideal timing, and consider team performance metrics to prevent disruptions during critical tasks.
+Despite its insights, our study’s short duration and single-session focus may not fully capture the depth, intricacies, and dynamics of extended human-robot collaborations and how they are influenced by small talk.
+This work used a small sample size $(N=20)$ and only studied one-time interactions.
+Thus, prolonged effects are unclear, and future work is needed to investigate longer or multi-session interactions.
+Additionally, the study did not consider any correlations between participants’ cultural nuances and their small talk engagements; this is vital as separate cultures perceive and value small talk differently (Cui, 2015; Rings, 1994; Bickmore and Cassell, 2005).
+The study also did not accurately evaluate its intended target demographic, as participants were recruited from a local university campus and surrounding areas instead of focusing on industrial or warehouse workers.
+Moreover, it was conducted in a lab setting with a low-stakes, simulated warehouse task.
+While this work examines the use of small talk in human-robot collaboration, future studies should focus on quantitatively evaluating the benefits of small talk for rapport building, how the user’s personality, cultural background, and task context may affect their small talk interaction, and the role of small talk with cognitively demanding user tasks.
+Ensuring the appropriateness of content, frequency, timing, and context are essential in designing conversational aspects of robot small talk in future research.
+For instance, a robot conversational system could be sensitive to user preferences on topics, adapt to individual users for ideal timing, and consider performance metrics to prevent disruptions during critical tasks; it is necessary to study the effects of different aspects of a robot’s conversational characteristics.
 Lastly, future research should explore methods to align a robot’s conversational capabilities with its functional abilities.
 It is important to keep users apprised of the robot’s true abilities, possibly using small talk as an informative medium.
-Nevertheless, the strategic introduction of small talk can be instrumental in building and strengthening rapport between users and task-oriented robots in collaborative scenarios.
+6. Conclusion
+This study explored user engagement with robots that initiate small talk in a simulated industrial collaborative setting.
+Our findings revealed a general willingness among users to engage in small talk with robots.
+Although users are receptive to robotic systems with small talk capabilities, which suggests potential enhancements to interaction experiences, we lack concrete evidence of the measurable benefits of robot-initiated small talk.
+Important interaction characteristics were identified, emphasizing the need for flexibility in the robot systems to adapt small talk’s frequency, complexity, and content to suit various users and tasks.
+However, introducing small talk could potentially distract users and reduce task focus and productivity, particularly during critical tasks.
+Therefore, limiting or avoiding small talk in such situations might be prudent.
+While social chat with robots is not new, integrating robot small talk in physical, collaborative settings with humans is a newer concept.
+Given the increasing prevalence of such robots in real-world settings and the existing interest in social interaction among industrial workers, we recommend further research to implement and quantify the benefits of social conversations between humans and robots in these environments.
 Acknowledgements.
@@
 Appendix
-Table 3. Overview of Codes and Their Definitions for the Data Analysis
+Table 6. Overview of Codes and Their Definitions for the Data Analysis: Part A
+Robot’s Verbal Phrases: Task-Oriented
+Name Definition
+Confirmation Robot confirms it processed user request
+Reminder Robot gives a reminder message about the task
+Robot’s Verbal Phrases: Small-Talk
+Name Definition
+Question Robot initiates a question to the user
+Statement Robot makes a small-talk response or comment
+Participant’s Verbal Phrases: Task-Oriented
+Name Definition
+Request User requests a yellow or green pipe, or reports a faulty joint
+Reminder User responses to the robot’s reminder message
+Confirmation User responses to the robot’s confirmation message
+System Announcement User responses to the negative system announcement message
+Participant’s Verbal Phrases: Small Talk
+Name Definition
+Question User initiates a question to the robot
+Repeat User asks the robot to repeat or clarify what it just said
+Short Non-question small talk statement ¡= 5 words
+Medium Non-question small talk statement ¿ 5 words
+Long Multi-sentence non-question small talk statements
+Table 7. Overview of Codes and Their Definitions for the Data Analysis: Part B
+Motives for Initiating Questions
+Name Definition
+Curiosity Individual describes a level of curiosity
+Felt Human-like Individual makes connections to human-likeness
+Social Convention Individual describes factors relating to social conventions
+Non-Verbal Reactions to Negative Feedback
+Name Definition
+Verbal Response An audible or verbal phrase spoken by the participant
+Non-Verbal Response Behaviors such as smiling, laughing, eye-rolls, head nods, etc.
+Blame Attribution
+Name Definition
+Self Selects themselves as a potential cause for the feedback announcement
+Robot Selects the robot as a potential cause for the feedback announcement
+Team Selects the team as a potential cause for the feedback announcement
+Nobody Believes neither they nor the robot caused the feedback announcement
 Response to First Robot Question: Conversation Reciprocation
 Name Definition
-Answer
-Response (Yes) and answered robot’s question
-Refuse
-Response (Yes) but did not answer the robot’s question
-No engagement
-Participant did not answer the robot’s question
+Answer Response (Yes) and answered robot’s question
+Refuse Response (Yes) but did not answer the robot’s question
+No engagement Participant did not answer the robot’s question
 Response to First Robot Question: Impact on Task
 Name Definition
-Minimal (glance)
-Looks at the robot but no task disruption
-Freeze
-Stops current action (minimum of 2 seconds) and looks at the robot before continuing the task
-None
-Did not stop or look at the robot
+Minimal (glance) Looks at the robot but no task disruption
+Freeze Stops current action (min. 2 sec) and looks at robot before continuing
+None Did not stop or look at the robot
 Response to First Robot Question: Affective Reactions
 Name Definition
-Pleasant
-Individual exhibits a smile
-Surprise
-Individual exhibits behaviors such as a chin raise, lower lip raise, brow lowering, etc.
-Indifference
-No reaction
-Non-Verbal Reactions to Negative Feedback
-Name Definition
-Verbal Response
-An audible or verbal phrase spoken by the participant
-Non-Verbal Response
-Behaviors such as smiling, laughing, eye-rolls, head nods, arm movements, etc.
-Blame Attribution
-Name Definition
-Self
-Selects themselves as a potential cause for the negative feedback announcement
-Robot
-Selects the robot as a potential cause for the negative feedback announcement
-Team
-Selects the team as a potential cause for the negative feedback announcement
-Nobody
-Believes neither they nor the robot caused the negative feedback announcement
-Motives for Initiating Questions
-Name Definition
-Curiosity
-Individual describes a level of curiosity
-Felt Human-like
-Individual makes connections to human-likeness
-Social Convention
-Individual describes factors relating to social conventions
-l6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">
+Pleasant Individual exhibits a smile
+Surprise Individual exhibits behaviors such as a chin raise, lower lip raise, brow lowering, etc.
+Indifference No reaction